### v1

# Nash Equilibria in Games with Playerwise Concave Coupling Constraints: Existence and Computation

Abstract
We study the existence and computation of Nash equilibria in continuous static games where the players’ admissible strategies are subject to shared coupling constraints, i.e., constraints that depend on their joint strategies. Specifically, we focus on a class of games characterized by playerwise concave utilities and playerwise concave constraints. Prior results on the existence of Nash equilibria are not applicable to this class, as they rely on strong assumptions such as joint convexity of the feasible set. By leveraging topological fixed point theory and novel structural insights into the contractibility of feasible sets under playerwise concave constraints, we give an existence proof for Nash equilibria under weaker conditions. Having established existence, we then focus on the computation of Nash equilibria via independent gradient methods under the additional assumption that the utilities admit a potential function. To account for the possibly nonconvex feasible region, we employ a log barrier regularized gradient ascent with adaptive stepsizes. Starting from an initial feasible strategy profile and under exact gradient feedback, the proposed method converges to an $\epsilon$-approximate constrained Nash equilibrium within $\mathcal{O}(\epsilon^{-3})$ iterations.
1 INTRODUCTION
Optimization problems involving multiple self-interested agents with coupled objectives frequently arise in applications across robotics, computer science, and economics. Game theory provides a natural framework for modelling such interactions. In many interactive tasks, players are restricted to optimizing over a set of admissible strategies that are influenced by the choices of others. For instance, players may be subject to satisfying safety requirements such as collision avoidance in multi-robot control (Gu et al.,, 2023), or operating within a shared budget, e.g. limited bandwidth in communication networks (Lasaulce and Tembine,, 2011). This more general game setting is captured by introducing coupling constraints, i.e., constraints that depend on the players’ joint strategy profile. In economic modeling, such constraint games have been applied to analyze equilibria under resource limitations and market constraints (Arrow and Debreu,, 1954).
In the following, we distinguish between games with non-shared and shared coupling constraints. In the former, more general setting, each player may face distinct coupling constraints—a framework introduced by Debreu, (1952). Therein, constrained Nash equilibria, referred to as social equilibria, are shown to exist in games with compact, convex strategy spaces and playerwise quasi-concave utilities, provided that each player has a nonempty set of feasible response strategies, irrespective of the strategies chosen by others. This assumption is strong: in many practical scenarios, such as under shared safety constraints, violations caused by a subset of players may leave others unable to restore feasibility. However, without any additional assumptions, we give a simple example of a game with linear utilities and non-shared linear coupling constraints for which no constrained Nash equilibrium exists (see Appendix B).
This motivates studying the latter case: shared coupling constraints, where the set of admissible joint strategies is identical across players. In this setting, Rosen, (1965) showed that a constrained Nash equilibrium exists under the assumption of joint convexity of the feasible region. However, even bilinear coupling constraints in two-player games are generally nonconvex. Playerwise concavity, a relaxation of joint concavity, is a common structural assumption on utilities that ensures the existence of Nash equilibria in games without coupling constraints (Nikaidô and Isoda,, 1955). For games with coupling constraints where the feasible region is defined by shared, playerwise concave constraint functions, past work has not addressed the existence of Nash equilibria.
On the computational side, it is important to acknowledge two fundamental limitations: (a) Even for shared bilinear constraints, finding a feasible strategy is NP-hard (Witsenhausen,, 1986). And (b), even in games without coupling constraints, for a general-sum utility structure, finding a Nash equilibrium is PPAD-hard (Daskalakis et al.,, 2009).
To circumvent the intractability associated with (a), given an initial feasible strategy, one can ask whether there exist optimization methods under which players converge to a constrained Nash equilibrium. Under this initial feasibility assumption, recent work has provided positive answers in the context of constrained Markov potential games (Alatur et al.,, 2024; Jordan et al.,, 2024). Despite the limits posed by (b), convergent learning methods have been studied for several subclasses where Nash equilibria can be efficiently computed, including zero-sum games (Cesa-Bianchi and Lugosi,, 2006) and potential games (Heliou et al.,, 2017; Anagnostides et al.,, 2022).
To further position our work within the literature, we note that in the context of the variational inequality framework, a substantial body of work addresses games with coupling constraints. Recent results (Jordan et al.,, 2023; Tatarenko and Kamgarpour,, 2024) provide algorithms with convergence to constrained Nash equilibria for the class of monotone games with jointly linear constraints. Thus, their assumptions on utilities as well as constraints are stronger than our playerwise concavity. Coupling constraints that induce nonconvex feasible regions are considered in Scarabaggio et al., (2024), who do not assume playerwise concavity of constraints and characterize a weaker notion of local equilibria.
In summary, to the best of our knowledge, the class of continuous constrained games with playerwise concave objectives and shared, playerwise convex coupling constraints has not been explicitly studied. Applying prior results to this setting requires introducing restrictive assumptions. In this work, we focus on the existence and computation of equilibria in this class of games.
In the first part, we investigate the existence of constrained Nash equilibria. Our approach generalizes the result of Rosen, (1965) by relaxing the assumption on the feasible region from joint convexity to playerwise convexity. We build on tools from topological fixed point theory, specifically the works of Eilenberg and Montgomery, (1946) and Begle, (1950), which extend fixed point results from convex to more general, contractible domains. Analyzing the topological structure of the feasible region for the considered game class is a key technical contribution of our work.
In the second part of the paper, we address the computational problem. Our focus is on independent learning algorithms, where players simultaneously update their strategies using only gradient feedback on their own utility and the shared constraints, without requiring additional communication. We draw inspiration from log barrier regularized interior point methods. Originally proposed by Frisch, (1955), this approach has been extensively studied in the context of constrained optimization problems (Megiddo,, 1989; Hinder and Ye,, 2024), but has remained unexplored in the context of games. By carefully selecting adaptive learning rates, we guarantee convergence to approximate constrained Nash equilibria in a subclass of playerwise concave games with shared coupling constraints, namely when utilities admit a potential function.
Contributions
In summary, we consider continuous games with playerwise concave utilities and shared, playerwise concave coupling constraints. Our contributions are as follows:
  • We establish the existence of constrained Nash equilibria without relying on restrictive assumptions made in prior work. Our approach provides new insights into the topological structure of the feasible region.
  • We propose an independent log barrier regularized gradient method. Starting from an initial feasible strategy, players update their strategies based on exact gradient feedback for utility and constraint functions, using adaptive learning rates. For constrained potential games, we prove best-iterate convergence to a constrained $\epsilon$-approximate Nash equilibrium strategy at an $\mathcal{O}(\epsilon^{-3})$ rate.
  • We empirically demonstrate convergence of the proposed algorithm in an identical interest game with nonconvex feasible region, and in a network routing game with coupled link capacity constraints.
2 PRELIMINARIES
Notation
For a finite set $X$, we denote the probability simplex over $X$ by $\Delta(X)$. For $n\in\mathbb{N}$, we use the notation $n]\coloneqq\left\{1,\dots,n\right\}$. For $x\in\mathbb{R}^{n}$, let $\lVert x\rVert$ denote the $\ell_{2}$-norm.
Game setting
In this paper, we study static games with continuous strategy spaces. Let $[m]$ be the set of players. For each $i\in[m]$, $\mathcal{X}_{i}\subset\mathbb{R}^{d_{i}}$ is a compact convex set of available strategies. Joint strategy profiles are denoted by $x=(x_{1},\dots,x_{m})\in\mathcal{X}\coloneqq\times_{i\in[m]}\mathcal{X}_{i}$. Upon playing strategy $x\in\mathcal{X}$, each player $i$ receives utility $u_{i}(x)$ where $u_{i}:\mathcal{X}\to\mathbb{R}$ is continuous. Shared coupling constraints are introduced through the set of functions and thresholds $\{(c_{j},\alpha_{j})\}_{j\in[b]}$ where for each $j\in[b]$, $\alpha_{j}\in\mathbb{R}$, and $c_{j}:\mathcal{X}\to\mathbb{R}$ is a continuous function. Players are assumed to be utility maximisers subject to lower bounds on the constraint functions. The set of strategy profiles considered feasible is given by
$\mathcal{C}\coloneqq\left\{x\in\mathcal{X}\;\big|\;\forall j\in[b],\;c_{j}(x)\geq\alpha_{j}\right\}.$
For joint strategies of all players other than $i$, we use the notation $x_{-i}\coloneqq(x_{1},\dots,x_{i-1},x_{i+1},\dots,x_{m})\in\mathcal{X}_{-i}\coloneqq\times_{j\in[m]\setminus\{i\}}\mathcal{X}_{j}$. For $i\in[m]$ and $x_{-i}\in\mathcal{X}_{-i}$, let $\mathcal{C}_{i}(x_{-i})\coloneqq\{x_{i}\in\mathcal{X}_{i}\mid(x_{i},x_{-i})\in\mathcal{C}\}$.
Nash equilibria
For games without coupling constraints, a Nash equilibrium is a strategy profile $x\in\mathcal{X}$ such that for all $i\in[m]$ and playerwise deviations $x^{\prime}_{i}\in\mathcal{X}_{i}$, we have $u_{i}(x^{\prime}_{i},x_{-i})\leq u_{i}(x)$. In our constrained setting, $x\in\mathcal{X}$ is called constrained Nash equilibrium if $x\in\mathcal{C}$ and for all $i\in[m]$ and $x^{\prime}_{i}\in\mathcal{C}_{i}(x_{-i})$, it holds that $u_{i}(x^{\prime}_{i},x_{-i})\leq u_{i}(x)$.
Remark 1.
The term constrained game has been used in the literature with varying meanings. In some cases, it has been used to distinguish games over convex compact strategy spaces from unconstrained games over $\mathbb{R}^{n}$. Our notion of constrained game is more general, as it additionally allows for shared coupling constraints.
Playerwise concavity of utilities is commonly assumed for proving existence of unconstrained Nash equilibria. We additionally make an equivalent playerwise concavity assumption on the constraints.
Assumption 1.
For all $i\in[m]$ and $x_{-i}\in\mathcal{X}_{-i}$, suppose $x_{i}\mapsto u_{i}(x_{i},x_{-i})$ and $x_{i}\mapsto c_{j}(x_{i},x_{-i})$ are concave for all $j\in[b]$.
3 EXISTENCE OF CONSTRAINED NASH EQUILIBRIA
In this section, we establish our existence result for constrained Nash equilibria. We first discuss assumptions made in prior work, then give an overview of our approach.
3.1 Discussion of assumptions
In the following, we motivate our assumptions using two examples of constrained games. Example 1 serves to illustrate that Assumption 1 is not sufficient for the existence of constrained Nash equilibria as it allows for feasible sets with nonsmooth boundary. We then introduce Assumption 2 to prevent the occurence of such degeneracies. Finally, we present Example 2, which satisfies both Assumption 1 and Assumption 2. At the same time, Example 2 illustrates that these assumptions are weaker than the conditions imposed by earlier existence results in the literature.
Figure 1: Illustration of the feasible sets $\mathcal{C}^{(1)}$ and $\mathcal{C}^{(2)}$ described in Example 1 and Example 2, respectively.
Example 1.
Consider a two-player game with continuous strategy spaces $\mathcal{X}_{1}=\mathcal{X}_{2}=[0,1]$ and feasible set $\mathcal{C}^{(1)}$ defined by a constraint $c_{1}(x_{1},x_{2})\coloneqq(x_{1}-\frac{1}{2})(x_{2}-\frac{1}{2})$ with threshold $\alpha_{1}=0$; illustrated on the left-hand side of Figure 1. Moreover, define the players’ utilities as $u_{1}(x_{1},x_{2})=u_{2}(x_{1},x_{2})\coloneqq x_{1}(1-x_{2})$. Observe that no strategy $(x_{1},x_{2})\in\mathcal{C}^{(1)}$ with $x_{1}<\frac{1}{2}$ can be a constrained Nash equilibrium, as $u_{1}$ would increase if player 1 would switch to $x_{1}^{\prime}=\frac{1}{2}$ and since $(x_{1}^{\prime},x_{2})\in\mathcal{C}^{(1)}$. By an analogous argument, no $(x_{1},x_{2})\in\mathcal{C}^{(1)}$ with $x_{2}>\frac{1}{2}$ can be a constrained Nash equilibrium. The only point not covered by above two cases is $(\frac{1}{2},\frac{1}{2})$ which however is not a Nash equilibrium either.
We note that in Example 1, the gradient of the constraint $c_{1}$ vanishes at a point that lies on the boundary of $\mathcal{C}^{(1)}$, namely $(\frac{1}{2},\frac{1}{2})$. In order to prove our existence result, we need to rule out such degeneracies. We therefore require the following playerwise variant of the Mangasarian-Fromovitz constraint qualification condition (Mangasarian and Fromovitz,, 1967); a regularity assumption that guarantees a uniform descent direction for all constraints at points on the boundary of $\mathcal{C}$. In our proof, this is used argue about continuity of best response mappings.
Assumption 2.
Let $\mathcal{J}(x)\coloneqq\left\{j\in[b]\mid c_{j}(x)=\alpha_{j}\right\}$ be the set of active constraints for a feasible strategy $x\in\mathcal{C}$. We assume that for all $x\in\mathcal{C}$ and $i\in[m]$, there exists $\delta_{i}\in\mathbb{R}^{d_{i}}$ such that for all $j\in\mathcal{J}(x)$, $\langle\delta_{i},\nabla_{x_{i}}c_{j}(x)\rangle>0$.
Finally, we present an example of a constrained game that satisfies Assumptions 1 and 2. Moreover, its feasible region $\mathcal{C}^{(2)}$ does not satisfy the stronger conditions required by previous existence results.
Example 2.
Let again $\mathcal{X}_{1}=\mathcal{X}_{2}=[0,1]$ and consider the feasible set $\mathcal{C}^{(2)}$ defined by two constraints, $c_{1}(x_{1},x_{2})\coloneqq(x_{1}-\frac{1}{2})(x_{2}-\frac{1}{2})-\frac{1}{111}$ and $c_{2}(x_{1},x_{2})\coloneqq\frac{1}{11}-(x_{1}-\frac{1}{2})(x_{2}-\frac{1}{2})$, with thresholds $\alpha_{1}=\alpha_{2}=0$, as shown on the right-hand side of Figure 2. The utility structure is irrelevant for this example. The resulting feasible set $\mathcal{C}^{(2)}$ is illustrated as the blue region on the right-hand side of Figure 1. The constraints satisfy Assumption 1, as fixing $x_{1}$ yields a linear function in $x_{2}$, and vice versa. Moreover, Assumption 2 holds, as $|\mathcal{J}(x)|=1$ for all $x\in\mathcal{C}^{(2)}$ and neither $c_{1}$ nor $c_{2}$ has any critical points on the boundary of $\mathcal{C}^{(2)}$. One can also observe that $\mathcal{C}^{(2)}$ is not convex, hence Rosen, (1965)’s result is not applicable. Moreover, for any $x\in\mathcal{X}$ with $x_{1}=0.5$, we have $\mathcal{C}^{(2)}_{2}(x_{1})=\emptyset$. Hence, results relying on the nonemptiness of feasible responses such as Debreu, (1952) are not applicable either.
Feasible regions similar to the one in Example 2 naturally arise in matrix games, where nonconvexity stems from joint probability products, and the nonemptiness condition fails when satisfying the constraint requires some coordination among players to choose certain feasible joint actions, see e.g. the applications in Section 5.
3.2 Existence result
We now state our existence theorem. Note that clearly no Nash equilibrium can exist when $\mathcal{C}=\emptyset$. {restatable}theoremthmexistence Let Assumptions 1 and 2 hold and suppose $\mathcal{C}\not$$\emptyset$. Then there exists a constrained Nash equilibrium $x^{\star}\in\mathcal{X}$. The remainder of this section is devoted to giving an overview of our proof approach and introducing the relevant topological background.
Fixed point theory
Existence proofs for Nash equilibria typically proceed by constructing a best response map to which a fixed point theorem is applied. Classical results such as Brouwer’s or Kakutani’s theorem require the domain of this map to be convex. In order to establish the existence of a fixed point for a best response map defined over our nonconvex $\mathcal{C}$, we rely on a result that goes back to Eilenberg and Montgomery, (1946), and Begle, (1950), which relaxes the convexity requirement to the weaker topological notion of contractibility. Below, we define contractibility and state a version of Begle, (1950)’s result applied to finite-dimensional Euclidean space. For examples of contractible sets, and a proof of Fact 3.2, see Appendix C. {restatable}definitiondefcontractibility Let $S\subseteq\mathbb{R}^{n}$. We say $S$ is contractible if there exists $s_{0}\in S$, and a continuous map $H:[0,1]\times S\to S$ such that for all $s\in S$, it holds that $H(0,s)=s$ and $H(1,s)=s_{0}$. {restatable}factfactfp Let $X\subset\mathbb{R}^{n}$ be compact, contractible, and $\phi:X\to X$ continuous. Then there exists $x\in X$ such that $x=\phi(x)$, that is, $x$ is a fixed point of $\phi$.
As seen in Example 2, the feasible region may be disconnected; thus, it cannot be contractible. Therefore, we define a best response function only over a connected component of $\mathcal{C}$, to which we then apply Fact 3.2. We conclude this section by sketching the proof of Theorem 3.2. The formal proof is given in Appendix D.
Figure 2: Proof sketch for contractability of $\mathcal{C}^{\prime}$.
Proof overview
Let $\mathcal{C}^{\prime}\subseteq\mathcal{C}$ be any connected component of $\mathcal{C}$, which exists as $\mathcal{C}\not$$\emptyset$. For $x_{-i}\in\mathcal{X}_{-i}$, we use the notation $\mathcal{C}^{\prime}_{i}(x_{-i})\coloneqq\{x_{i}\in\mathcal{X}_{i}\mid(x_{i},x_{-i})\in\mathcal{C}^{\prime}\}$ and $\operatorname{supp}(\mathcal{C}_{i}^{\prime})\coloneqq\left\{x_{-i}\in\mathcal{X}_{-i}\mid\mathcal{C}^{\prime}_{i}(x_{-i})\not$$\emptyset\right\}$. The proof is divided into two parts:
  1. 1)
    Contractibility of $\mathcal{C}^{\prime}$: We construct a mapping $H:[0,1]\times\mathcal{C}^{\prime}\to\mathcal{C}^{\prime}$ by sequentially contracting the feasible region along the dimensions of the players. For this overview, we focus on the two-player case which is illustrated in Figure 2.
    For $t\in[0,\frac{1}{2}]$ and $x=(x_{1},x_{2})\in\mathcal{C}^{\prime}$, we let $t\mapsto H(t,x)$ linearly interpolate between $x$ and $(x_{1},\overline{c}(\mathcal{C}^{\prime}_{2}(x_{1})))$ where for compact $S\subset\mathbb{R}^{n}$, $\overline{c}(S)$ denotes the centroid of $S$. This is depicted by the green arrows in Figure 2.
    For $t\in[\frac{1}{2},1]$ and $x=(x_{1},x_{2})\in\mathcal{C}^{\prime}$, we let $H(t,x)$ be the point $(x_{1}(t),\mathcal{C}^{\prime}_{2}(x_{1}(t)))$ where $x_{1}(t)$ linearly interpolates between $x_{1}$ and $\overline{c}(\operatorname{supp}(\mathcal{C}^{\prime}_{1}))$. This is depicted by the red arrows in Figure 2.
    In order to satisfy Definition 3.2, $H$ needs to be (a) well-defined, such that for all $t\in[0,1]$ and $x\in\mathcal{C}^{\prime}$, we have $H(t,x)\in\mathcal{C}^{\prime}$, (b) continuous in $(t,x)$, and (c) must contract $\mathcal{C}^{\prime}$ to some point $\hat{x}\in\mathcal{C}^{\prime}$.
    Property (a) is shown via convexity of the slices $\mathcal{C}^{\prime}_{i}(x_{-i})$ which follows directly from the playerwise concavity on the constraints. Regarding (b), we observe that $\overline{c}(\cdot)$ and $\mathcal{C}_{i}^{\prime}(\cdot)$, from which we compose $H$, are both continuous. In particular, continuity of the set-valued mapping $\mathcal{C}_{i}^{\prime}(\cdot)$ follows from standard results in parametric optimization, which apply under the MFCQ condition (Assumption 2). For (c), by construction, for all $x\in\mathcal{C}^{\prime}$, $H(0,x)=x$ and $H(1,x)=\overline{c}(\operatorname{supp}\mathcal{C}_{1}^{\prime})$.
  1. 2)
    Fixed point argument: Our goal is to define a best response function $\Psi:\mathcal{C}^{\prime}\to\mathcal{C}^{\prime}$, to which Fact 3.2 can be applied. We leverage the playerwise concave constraint structure by constructing $\Psi$ in a playerwise sequential manner. For each $i\in[m]$, suppose $u_{i}$ is strongly concave (we will justify why this is without loss of generality). Then $\Psi_{i}:\mathcal{C}^{\prime}\to\mathcal{C}^{\prime}$ with
    $\Psi_{i}(x)\coloneqq\Big(\operatorname*{arg\,max}_{x_{i}^{\prime}\in\mathcal{C}^{\prime}_{i}(x_{-i})}u_{i}(x_{i}^{\prime},x_{-i}),\;x_{-i}\Big).$
    is well-defined since the maximizer is unique. Moreover, by a variant of Berge’s maximum theorem for strongly concave objective functions over convex constraints (see Theorem 1 in Appendix C), $\Psi_{i}$ is continuous. We then apply Fact 3.2 to the composition $\Psi\coloneqq\Psi_{1}\circ\dots\circ\Psi_{m}$. Using the playerwise concavity of constraints, it can be argued that any fixed point $x^{\star}\in\mathcal{C}^{\prime}$ with $x^{\star}=\Psi(x^{\star})$ is a constrained Nash equilibrium.
    Why is assuming strict concavity of $u_{i}$ without loss of generality? For $\epsilon>0$, consider the perturbed game with playerwise strongly concave utilities $u_{i}^{\epsilon}(x)\coloneqq u_{i}(x)-\epsilon\lVert x_{i}\rVert^{2}$. By a limit argument, we can show that for a sequence of such games with $\epsilon\to 0$, any limit point of the respective constrained Nash equilibria (just shown to exist) is a constrained Nash equilibrium of the original unperturbed game.
    We point out that, different from other existence proofs (Debreu,, 1952; Rosen,, 1965), it is important for the best response maps $\Psi_{i}$ to be single-valued: Otherwise it is not clear how to ensure $\Psi$ to be convex-valued (or contractible-valued), which is necessary for fixed point theorems to hold.
Having established existence, we next investigate settings in which such equilibria can be computed or learned in an efficient manner.
4 LEARNING CONSTRAINED NASH EQUILIBRIA
In this section, we turn to the problem of finding approximate constrained Nash equilibria in a deterministic, gradient-based, independent learning setting.
First, we introduce an approximate notion of constrained Nash equilibrium that serves as the solution concept for our algorithms. Note that the case $\epsilon=0$ recovers the exact notion from Section 2.
Definition 1.
For $\epsilon>0$, $x\in\mathcal{X}$ is a constrained $\epsilon$-approximate Nash equilibrium if $x\in\mathcal{C}$ and for all $i\in[m]$ and $x^{\prime}_{i}\in\mathcal{C}_{i}(x_{-i})$, it holds that $u_{i}(x^{\prime}_{i},x_{-i})\leq u_{i}(x)+\epsilon$.
While Assumptions 1 and 2 suffice for existence, without additional assumptions on the game structure, the problem of finding $\epsilon$-approximate Nash equilibria is known to be PPAD-hard even for unconstrained normal-form games (Daskalakis et al.,, 2009; Rubinstein,, 2015). Therefore, in Section 4.1, we focus on the commonly studied tractable sublass of potential games. Before specializing to this case, we introduce the considered independent learning protocol, discuss challenges and additional assumptions, and outline our general approach.
Independent learning protocol
We assume players to interact with the game for $T$ rounds. In each round $t$, for $0\leq t\leq T-1$, player $i$ chooses some strategy $\smash{x_{i}^{(t)}\in\mathcal{X}^{i}}$ and receives first-order feedback for the utility, i.e. $\nabla_{x_{i}}u_{i}(x^{(t)})$, as well as zeroth- and first-order feedback for each constraint $j\in[b]$, i.e., $c_{j}(x^{(t)})$ and $\nabla_{x_{i}}c_{j}(x^{(t)})$. Players do not observe other players’ strategies or utility feedback. Moreover, besides the game interaction, no communication or coordination (e.g., shared randomness) among players is allowed.
Smoothness and Lipschitz continuity of utility functions are standard assumptions in gradient-based learning, see e.g. Mertikopoulos et al., (2024); we extend these regularity conditions to the constraints.
Assumption 3.
There exist constants $L$ and $M$ such that the functions $u_{i}$ for $i\in[m]$ and $c_{j}$ for $j\in[b]$ are all differentiable, $L$-Lipschitz, and $M$-smooth in the joint strategies $x\in\mathcal{X}$.
Moreover, we make the following assumption on strict feasibility of the initial strategy.
Assumption 4.
There exist $\gamma>0$ and $x^{(0)}\in\mathcal{X}$ such that $\min_{j\in[b]}\{c_{j}(x^{(0)})-\alpha_{j}\}\geq\gamma$. For all $i\in[m]$, we assume $x^{(0)}_{i}$ is initially known to player $i$.
Note that without assuming initial feasibility, finding a feasible strategy in a two-player game with bilinear constraints is as hard as finding a global maximizer of a bilinear function. The latter problem is known to be NP-hard  (Witsenhausen,, 1986) even in a centralized setting.
Furthermore, preserving the initial feasibility across subsequent iterates motivates the use of interior point methods. Our method is inspired by log barrier regularized gradient descent, an interior point method that has recently been studied in the context of nonconvex optimization under nonconvex constraints (Hinder and Ye,, 2024; Usmanova et al.,, 2024).
Log barrier regularization
First, we define the set $\mathcal{C}^{\circ}\coloneqq\left\{x\in\mathcal{X}\;\big|\;\forall j\in[b],\;c_{j}(x)>\alpha_{j}\right\}$ of strictly feasible strategies. For each player $i\in[m]$ and $x\in\mathcal{C}^{\circ}$, let
$B_{i}^{\eta}(x)\coloneqq u_{i}(x)+\eta\sum_{j\in[b]}\log(c_{j}(x)-\alpha_{j})$
where $\eta>0$ is a parameter controlling the regularization. Observe that for each constraint $j$, the log-term serves as a barrier that tends to -$\infty$ as $c_{j}(z)$ approaches the threshold $\alpha_{j}$ for any $j\in[b]$. Then, we propose the following independent log barrier regularized gradient ascent: Suppose in iteration $t$ for $0\leq t\leq T-1$, each player $i$ simultaneously performs the projected regularized gradient update
$x_{i}^{(t+1)}=\mathcal{P}_{\mathcal{X}_{i}}\left[x^{(t)}+\gamma^{(t)}\nabla_{x_{i}}B^{\eta}_{i}(x)\right]$ (1)
for some stepsize $\gamma^{(t)}>0$, where $\nabla_{x_{i}}B_{i}^{\eta}(x)=\nabla_{x_{i}}u_{i}(x)+\eta\sum_{j\in[b]}\frac{\nabla_{x_{i}}c_{j}(x)}{c_{j}(x)-\alpha_{j}}$, and $\mathcal{P}_{\mathcal{X}_{i}}[\cdot]$ denotes projection onto $\mathcal{X}_{i}$. Our goal in the following is to design $\smash{\gamma^{(t)}}$ and $\eta$ to guarantee feasibility of all joint iterates and convergence to a constrained Nash equilibrium.
Smoothness along trajectory
Even though we assume smoothness of utilities and constraints, $B^{\eta}_{i}(x)$ is not globally smooth due to unbounded growth of the log-term towards the boundary of $\mathcal{C}$. However, we show that $B^{\eta}_{i}(x)$ is smooth over the joint strategy trajectory of independent log barrier regularized gradient ascent when using suitable adaptive stepsizes $\smash{\gamma^{(t)}}$.
While for the existence result, the MFCQ condition stated in Assumption 2 is sufficient, our convergence proof requires an extended version of this condition. Namely, a uniform constraint ascent direction must exist for all feasible joint strategies $x\in\mathcal{C}$ near the boundary of $\mathcal{C}$, ensuring that log barrier gradients consistently steer the iterates away from the boundary.
Assumption 5.
For $\rho>0$, let $\mathcal{J}_{\rho}(x)\coloneqq\{j\in[b]\mid c_{j}(x)-\alpha_{j}\leq\rho\}$ be the set of $\rho$-approximately active constraints for strategy $x\in\mathcal{C}$. For some $\rho\in(0,\frac{\gamma}{2}]$, there exists $\ell>0$, such that for all $x\in\mathcal{X}$ and all $j\in\mathcal{J}_{\rho}(x)$, there exists a direction $\delta$ with $x+\delta\in\mathcal{X}$, $\lVert\delta\rVert\leq 1$, such that $\langle\delta,\nabla c_{j}(x)\rangle>\ell$.
Note that for $\rho=0$ and $\mathcal{X}=\mathbb{R}^{n}$, we recover the classic Mangasarian-Fromovitz constraint qualification (MFCQ) condition (Mangasarian and Fromovitz,, 1967).
4.1 Learning in Constrained Potential Games
We consider the class of games with coupling constraints that, when the constraints are omitted, reduce to a potential game, as introduced by Monderer and Shapley, (1996). For our purpose, it suffices to have a potential function defined over the feasible region.
Assumption 6.
Suppose there exists a potential function $\Phi:\mathcal{C}\to\mathbb{R}$ such that for all players $i\in[m]$, strategy profiles $x\in\mathcal{C}$, and playerwise deviations $x^{\prime}_{i}\in\mathcal{C}_{i}(x_{-i})$,
$\Phi(x)-\Phi(x^{\prime}_{i},x_{-i})=u_{i}(x)-u_{i}(x^{\prime}_{i},x_{-i}).$
Examples of such continuous potential games trivially include identical interest games, and, as shown in Example 3, mixed strategy extensions of finite action potential games. This class of potential games also arises in economics, e.g. in the context of the Cournot competition (Cournot,, 1838), in communication networks (Scutari et al.,, 2006), and in transportation (Altman and Wynter,, 2004).
Example 3.
Consider the unconstrained game with joint action space $\mathcal{A}\coloneqq\times_{i\in[m]}\mathcal{A}_{i}$ where each $\mathcal{A}_{i}$ is a finite set, and utilities $\widetilde{u}_{i}:\mathcal{A}_{i}\to[0,1]$ such that there exists a potential function $\phi:\mathcal{A}\to\mathbb{R}$ that satisfies $\phi(a)-\phi(a_{i}^{\prime},a_{-i})=\widetilde{u}_{i}(a)-\widetilde{u}_{i}(a_{i}^{\prime},a_{-i})$ for all $i\in[m]$, $a\in\mathcal{A}$, and $a^{\prime}_{i}\in\mathcal{A}_{i}$. Then, the mixed strategy extension of this game with strategy space $\mathcal{X}\coloneqq\times_{i\in[m]}\Delta(\mathcal{A}_{i})$ and utilities $u_{i}(x)\coloneqq\mathbb{E}_{a\sim x}[\widetilde{u}_{i}(a)]$ is a potential game with potential $\Phi(x)\coloneqq\mathbb{E}_{a\sim x}[\phi(a)]$. See Appendix F for a proof of this fact.
Next, we state our best-iterate convergence result for potential games with coupling constraints. An overview of the analysis is given below; for the full proof, we refer to Appendix F. {restatable}theoremthmpotential Let $\epsilon>0$ and set $\eta=\epsilon$. Let Assumptions 1, 3, 4, 6, and 5 hold with $\rho\geq\eta$. For stepsizes $\{\gamma^{(t)}\}_{t=0}^{T-1}$ chosen as in (2), suppose we perform $T=\mathcal{O}(\epsilon^{-3})$ rounds of simultaneous updates as in (1). Then, there exists $0\leq t\leq T$ such that $x^{(t)}$ is a constrained $\epsilon$-approximate Nash equilibrium.
Remark 2.
  1. 1.
    Compared to the corresponding result in unconstrained potential games, see Theorem B.6 of Anagnostides et al., (2022) which achieves a $\mathcal{O}(\epsilon^{-2})$ best-iterate rate, log barrier regularization comes at the cost of an additional $\mathcal{O}(\epsilon^{-1})$ factor. This is due to small stepsizes as iterates approach the boundary of the feasible region.
  2. 2.
    While we are not aware of any results on learning static potential games with playerwise concave coupling constraints, a related result exists for constrained Markov potential games under an independent learning protocol similar to ours (Jordan et al.,, 2024). Unlike our approach, it relies on a double-loop proximal-point-like algorithm and, in the exact-gradient setting, achieves best-iterate convergence to a constrained $\epsilon$-approximate Nash equilibrium a slower $\mathcal{O}(\epsilon^{-4})$ rate. Moreover, Jordan et al., (2024) requires a so-called uniform Slater’s condition which in practice may be harder to verify than our Assumption 5.
  3. 3.
    An appealing property of log barrier methods is the feasibility of all iterates, making them well-suited for problems where constraints model safety conditions that may not be violated during learning.
Proof overview
First, we observe that due to the identical interest structure of the log barrier regularization, $\Phi^{\eta}(x)\coloneqq\Phi(x)+\eta\sum_{j\in[b]}\log(c_{j}(x)-\alpha_{j})$ is a potential function for the regularized game. Therefore, partial gradients are aligned with the potential gradient, i.e., $\nabla_{x_{i}}B^{\eta}_{i}(x)=\nabla_{x_{i}}\Phi^{\eta}(x)$ for all $i\in[m]$ which allows us to equivalently analyze our independent updates as a centralized log barrier regularized gradient ascent on $\Phi^{\eta}$.
Next, we argue about smoothness of $\Phi^{\eta}$ along the gradient ascent trajectory. We say $x\in\mathcal{C}$ is $\beta$-feasible if $c_{j}(x)-\alpha_{j}\geq\beta$ for all $j\in[b]$. For $\beta$-feasible $x$, we give a Hessian bound $\lVert\nabla^{2}\Phi^{\eta}(x)\rVert\leq M^{\eta,\beta}_{\Phi}$ where $M^{\eta,\beta}_{\Phi}\coloneqq M\sqrt{m}+\eta Mb\beta^{-1}+\eta L^{2}b\beta^{-2}$. Moreover, we show that if an iterate $x^{(t)}$ is $\beta^{(t)}$-feasible, and the stepsize sequence is chosen adaptively such that for all $0\leq t\leq T-1$, it holds that $0<\gamma^{(t)}\leq\min_{j\in[b]}\frac{c_{j}(x^{(t)})-\alpha_{j}}{2mL^{2}}$, then all strategies on the line segment $x^{(t)},x^{(t+1)}]$ are $\beta^{(t)}/2$-feasible. Consequently, $\Phi^{\eta}$ is $M_{\Phi}^{\eta,\beta^{(t)}/2}$-smooth over $x^{(t)},x^{(t+1)}]$. Therefore, if we choose stepsizes
$\gamma^{(t)}\coloneqq\min\left\{\min_{j\in[b]}\frac{c_{j}(x^{(t)})-\alpha_{j}}{2mL^{2}},\frac{1}{M_{\Phi}^{\eta,\beta^{(t)}/2}}\right\},$ (2)
a standard inequality from nonconvex smooth optimization (Bubeck et al., (2015), Lemma 3.6) shows sufficient increase proportional to $\gamma^{(t)}$. Nevertheless, $M_{\Phi}^{\eta,\beta^{(t)}/2}$ may grow arbitrarily and lead to diminishing stepsizes. Leveraging Assumption 5, we give a lower bound on $\gamma^{(t)}$ based on the fact that the log barrier term dominates and steers iterates away from the boundary of $\mathcal{C}$ whenever the feasibility margin $\beta$ is too small.
As a result, for our choices of $\eta$ and $T$, we obtain best-iterate convergence at an $\mathcal{O}(\epsilon^{-3})$-rate towards an $\epsilon$-stationary point $x^{\star}$ of $\Phi^{\eta}$. We show that this implies a set of $\epsilon$-approximate KKT conditions to hold for the constrained optimization problem $\max_{x\in\mathcal{C}}\Phi^{\eta}(x)$ at $x^{\star}$, which in turn implies $x^{\star}$ to be an $\epsilon$-approximate constrained Nash equilibrium.
5 SIMULATIONS
To demonstrate the convergence of our method, we conduct numerical simulations on two continuous potential games with shared coupling constraints. The first is a simple two-player identical interest game played over a nonconvex feasible region. The second is a routing problem with link capacity constraints.
Identical interest game
Consider a two-player game where the strategy $x\in\mathcal{X}=[0,1]^{2}$ models each player’s contribution to a cooperative task for which utilities are given by
$u_{1}(x_{1},x_{2})$ $=u_{2}(x_{1},x_{2})$
$=\underbrace{(x_{1}+x_{2})}_{(a)}-a\underbrace{(x_{1}-x_{2})^{2}}_{(b)}-b\underbrace{(x_{1}^{2}+x_{2}^{2})}_{(c)}$
for some $a,b>0$. One may interpret term (a) as capturing the mutual benefit of the joint effort, term (b) as encouraging equal contribution, and term (c) as reflecting diminishing returns of individual high efforts. Additionally, we introduce a shared constraint by defining $c(x_{1},x_{2})\coloneqq 1-x_{1}x_{2}$ and threshold $\alpha=0.85$, representing a bound on the joint contribution.
Figure 3: Strategy trajectory in the utility landscape of the identical interest game; colors indicate the log barrier value; grey is infeasible.
Starting from an initial feasible point, we apply our log barrier regularized independent gradient ascent to the constrained potential game. Figure 3 visualizes the joint strategy trajectory within the utility landscape where colors indicate the log barrier value. The trajectory (in red) remains within the nonconvex feasible region throughout the iterations, aligning with our theoretical observations. In Figure 4, we plot the $\text{Nash-Gap}_{i}(x)\coloneqq\max_{x_{i}^{\prime}\in\mathcal{C}_{i}(x_{-i})}u_{i}(x_{i}^{\prime},x_{-i})-u_{i}(x)$ for each player $i$. This can be computed by solving the single-player constrained optimization problem where $x_{-i}$ remains fixed. The Nash gap serves as a measure of convergence, as for any $\epsilon>0$, if $\text{Nash-Gap}(x)\coloneqq\max_{i\in[m]}\text{Nash-Gap}_{i}(x)\leq\epsilon$, then $x$ is a constrained $\epsilon$-approximate Nash equilibrium. Hence, we observe convergence to a strategy near the boundary of $\mathcal{C}$ that forms an approximate constrained Nash equilibrium.
Figure 4: Nash gaps and constraint values over iterations $t$ for the identical interest game.
Network routing game
Moving beyond the identical interest case, we build upon an unconstrained distributed routing problem that has been shown to satisfy the potential structure (see Example 2.2, Tatarenko, (2017)). We extend this model by incorporating coupled capacity constraints. Specifically, 5 players, one corresponding to each color in Figure 5 aim to route some amount of flow through 5 paths in the network, each associated to the respective player ($R_{1}$: blue, $R_{2}$: red, $R_{3}$: brown, $R_{4}$: orange, $R_{5}$: green). For $i\in[5]$, we let $x_{i}\in[0,10]$ denote the flow on player $i$’s path $R_{i}$.
Figure 5: Routing game (Tatarenko, (2017), Ex. 2.2).
Utilities represent a players’ (linear) profit minus link costs, where the latter depends on the congestion of the used links. Concretely,
$u_{i}(x)\coloneqq px-\sum_{k:l_{k}\in R_{i}}P_{k}\Bigg(\sum_{j:l_{k}\in R_{j}}x_{j}\Bigg)$
where we set $p=42$ and $P_{1}(x)=5x^{2}+x$, $P_{2}(x)=4x^{2}+2x+1$, $P_{3}(x)=5x^{2}$, $P_{4}(x)=2x^{2}+x$, $P_{5}(x)=10x$. We additionally define shared capacity constraints on the links $l_{2}$ and $l_{5}$, namely
$c_{1}(x)\coloneqq\sum_{j:l_{2}\in R_{j}}x_{j},\quad\quad c_{2}(x)\coloneqq\sum_{j:l_{5}\in R_{j}}x_{j}$
with upper bounds $\alpha_{1}=2.7$ and $\alpha_{2}=7$. Note that by negating, these constraints can be formulated as lower bounds in order to match our problem definition. Moreover, utilities and constraints are playerwise concave due to linearity and convexity of the functions $P_{l}$ for $l\in[5]$, hence satisfying Assumption 1.
Starting from feasible flows $x_{i}=\frac{1}{2}$ for all $i\in[5]$, Figure 6 shows the convergence of log barrier regularized independent gradient ascent to a constrained Nash equilibrium of the routing game. Note that here $\text{Nash-Gap}(x^{(t)})\coloneqq\max_{i\in[5]}\text{Nash-Gap}_{i}(x^{(t)})$. We point out that, again, all iterates remain within the feasible region.
Figure 6: Nash gap and constraint values over iterations $t$ for the routing game.
6 CONCLUSION
In this paper, we established existence of Nash equilibria in continuous games with playerwise concave utilities and shared, playerwise concave constraints. We proposed an independent log barrier regularized gradient method and proved its convergence to constrained Nash equilibria in potential games with shared coupling constraints.
Our work suggests several directions for future work: Can equilibrium existence and computation be characterized when utilities or constraints lack playerwise concavity, as in Markov games? Do our results extend to settings with stochastic zeroth- or first-order feedback? And finally, one may ask whether our approach can be applied to learning weaker notions of equilibria in general-sum games, or constrained Nash equilibria in other tractable subclasses of playerwise concave games, such as zero-sum or monotone games.
Supplementary Materials
Appendix A Overview of Notation
Table 1 gives an overview of our notation. All notations are introduced before their first use as well.
Table 1: Overview of notation
$\Delta(X)$ $\coloneqq\quad$ probability simplex over finite set $X$
$[n]$ $\coloneqq\quad$ the set $\left\{1,2,\dots,n\right\}$ for $n\in\mathbb{N}$
$S^{\circ}$ $\coloneqq\quad$ interior of the set $S\subset\mathbb{R}^{n}$
$\mathcal{P}_{S}[x]$ $\coloneqq\quad$ projection of $x\in\mathbb{R}^{n}$ onto compact convex $S\subset\mathbb{R}^{n}$
$y]_{+}$ $\coloneqq\quad\max\left\{0,y\right\}$ for $y\in\mathbb{R}$
$\lVert x\rVert$ $\coloneqq\quad$ $\ell_{2}$-norm of $x\in\mathbb{R}^{n}$
$x,x^{\prime}]$ $\coloneqq\quad\left\{\lambda x+(1-\lambda)x^{\prime}\mid\lambda\in[0,1]\right\}\subset\mathbb{R}^{n}$ for $x,x^{\prime}\in\mathbb{R}^{n}$
$\operatorname{dist}(x,S)$ $\coloneqq\quad\inf_{x^{\prime}\in S}\lVert x-x^{\prime}\rVert$ for $S\subset\mathbb{R}^{n}$ and $x\in S$
$\operatorname*{conv}(S)$ $\coloneqq\quad$ convex hull of a set $S\subset\mathbb{R}^{n}$
$\mathcal{X}$ $\coloneqq\quad$ $\times_{i\in[m]}\mathcal{X}_{i}$, set of strategy profiles
$u_{i}(x)$ $\coloneqq\quad$ player $i$’s utility function $u_{i}:\mathcal{X}\to[0,1]$
$c_{j}(x)$ $\coloneqq\quad$ $j$-th shared cost function $c_{j}:\mathcal{X}\to[0,1]$
$\mathcal{C}$ $\coloneqq\quad$ set of feasible strategy profiles, $\mathcal{C}\subset\mathcal{X}$
$\mathcal{C}_{i}(x_{-i})$ $\coloneqq\quad\left\{x_{i}\in\mathcal{X}_{i}\mid(x_{i},x_{-i})\in\mathcal{C}\right\}$ for $x_{-i}\in\mathcal{X}_{-i}$
$\beta^{(t)}$ $\coloneqq\quad\min_{j\in[b]}\left\{c_{j}(x^{(t)})-\alpha_{j}\right\}$
$\gamma^{(t)}$ $\coloneqq\quad$ stepsize in iteration $t$ for (non-optimistic) log barrier gradient ascent
$B_{i}^{\eta}(x)$ $\coloneqq\quad u_{i}(x)+\eta\sum_{j\in[b]}\log(c_{j}(x)-\alpha_{j})$
$\Phi(x)$ $\coloneqq\quad$ potential function for $m$-player game in Section 4.1
$\Phi^{\eta}(x)$ $\coloneqq\quad\Phi(x)+\eta\sum_{j\in[b]}\log(c_{j}(x)-\alpha_{j})$; potential of regularized game
$M^{\eta,\beta}_{\Phi}$ $\coloneqq\quad$ Hessian norm bound for $\Phi^{\eta}(x)$ for $\beta$-feasible $x\in\mathcal{X}$
Appendix B Discussion on Constraint Structure
In this section, we provide further motivation for focusing our paper on the case of shared coupling constraints. We first define the setting with non-shared constraints. Then, we provide an example of a game for which no constrained Nash equilibrium exists.
Game setting
Consider the continuous static game setting with players $[m]$, joint strategy space $\mathcal{X}=\times_{i\in[m]}\mathcal{X}_{i}$, and utilities $u_{i}(x):\mathcal{X}\to\mathbb{R}$ for each $i\in[m]$, as before. Non-shared coupling constraints are introduced through the set of functions and thresholds $\left\{(c_{i,j},\alpha_{i,j})\right\}_{i\in[m],j\in[b_{i}]}$ where for each $i\in[m]$, $b_{i}$ is the number of constraints of player $i$. Moreover, for all $i\in[m]$ and $j\in[b_{i}]$, we have $\alpha_{i,j}\in\mathbb{R}$ and $c_{i,j}:\mathcal{X}\to\mathbb{R}$ continuous. The set of feasible strategy profiles is then given by
$\widetilde{\mathcal{C}}\coloneqq\left\{x\in\mathcal{X}\;\big|\;\forall i\in[m],j\in[b_{i}],c_{i,j}(x)\geq\alpha_{i,j}\right\}.$
For $i\in[m]$ and $x_{-i}\in\mathcal{X}_{-i}$, let the set of feasible responses of player $i$ to $x_{-i}$ be denoted by $\widetilde{\mathcal{C}}_{i}(x_{-i})\coloneqq\left\{x_{i}\in\mathcal{X}_{i}\mid\forall j\in[b_{i}],c_{i,j}(x)\geq\alpha_{i,j}\right\}$.
Nash equilibria
A constrained Nash equilibrium in the non-shared constraint setting is a strategy profile $x\in\mathcal{X}$ such that $x\in\widetilde{\mathcal{C}}$, and for all $i\in[m]$ and playerwise deviations $x_{i}^{\prime}\in\widetilde{\mathcal{C}}_{i}(x_{-i})$, it holds that $u_{i}(x_{i}^{\prime},x_{-i})\leq u_{i}(x)$.
We point out that in above definition, deviations of player $i$ only need to be feasible with respect to player $i$’s constraints; this is what distinguishes non-shared constraints from shared ones. Note that this equilibrium notion matches the one introduced by Debreu, (1952). The nonemptiness assumption of Debreu, (1952) then requires that for all $i\in[m]$ and $x_{-i}\in\mathcal{X}_{-i}$, we have $\mathcal{C}_{i}(x_{-i})\not$$\emptyset$. Under this assumption, Debreu, (1952) proves existence of constrained Nash equilibria.
Is it possible to extend Debreu, (1952)’s theorem and establish existence results based on playerwise concavity (our Assumption 1) in this class of games with non-shared constraints, without introducing further assumptions such as the above nonemptiness condition on the feasible response sets? We give a negative answer in the form of the following counterexample. Note that the counterexample even has jointly linear constraints and utilities.
Proposition 1.
There exists a two-player constrained game with non-shared playerwise concave constraints $c_{1,1}$ and $c_{2,1}$ that has no constrained Nash equilibrium.
Proof.
Let $\mathcal{X}_{1}=\mathcal{X}_{2}=[0,1]$ and define utilities and costs as follows:
$u_{1}(x_{1},x_{2})$ $=-x_{1},$
$u_{2}(x_{1},x_{2})$ $=-x_{2},$
$c_{1,1}(x_{1},x_{2})$ $=x_{2},$
$c_{2,1}(x_{1},x_{2})$ $=x_{1},$
with thresholds $\alpha_{1,1}=\alpha_{1,2}=0.5$. Assumption 1 on the playerwise concavity of utilities and constraints holds due to linearity of utility and constraint functions. The joint feasible region is
$\widetilde{\mathcal{C}}=\left\{x\in[0,1]^{2}\;\big|\;x_{1}\geq 0.5\land x_{2}\geq 0.5\right\},$
and for any $(x_{1},x_{2})\in[0,1]^{2}$, the feasible response sets are given by $\widetilde{\mathcal{C}}_{2}(x_{1})=\widetilde{\mathcal{C}}_{1}(x_{2})=[0.5,1]$. Observe that for any $x\in\widetilde{\mathcal{C}}$, player 1 can increase its utility by deviating to $x_{1}^{\prime}=0\in\widetilde{\mathcal{C}}_{1}(x_{2})$. Hence no constrained Nash equilibrium can exist. ∎
Note that Debreu, (1952)’s assumption prevents constructing such pathological cases: In above example, the nonemptiness condition is violated, as for any $\hat{x}\in[0,0.5)^{2}$, $\widetilde{\mathcal{C}}_{1}(\hat{x}_{2})=\widetilde{\mathcal{C}}_{2}(\hat{x}_{1})=\emptyset$. However, as seen in our Example 2, Debreu, (1952)’s nonemptiness assumption is strong and captures only a limited class of coupling constraints.
As we aim to avoid making further assumptions besides playerwise concavity, the above proposition motivates focusing on the case of shared constraints in this paper.
Appendix C Background for Section 3
In this section, we provide the background required for presenting a formal proof of our main existence result, Theorem 3.2.
[…]
Next, we provide a statement of Berge’s maximum theorem that is used in our existence proof. This particular variant considers maximization of strongly concave parameterized functions over a parameterized constraint set. For reference, see for instance the version stated in Sundaram, (1996), Theorem 9.17.
Theorem 1.
Let $X,\Theta\subseteq\mathbb{R}^{n}$, $f:X\times\Theta\to\mathbb{R}$, and $C:\Theta\to\mathcal{P}(X)$ be continuous, and for all $\theta\in\Theta$, let $C(\theta)$ be compact and nonempty. Suppose also that for all $\theta\in\Theta$, $f(\cdot,\theta)$ is strongly concave, and that $C(\theta)$ is convex. Then
  1. 1.
    $f^{\star}:\Theta\to\mathbb{R}$ with $f^{\star}(\theta)\coloneqq\max_{x\in C(\theta)}f(x,\theta)$ is well-defined and continuous, and
  2. 2.
    the function $C^{\star}:\Theta\to\mathcal{P}(X)$ defined by
    $\hskip-11.38092ptC^{\star}(\theta)$ $\coloneqq\arg\max_{x\in C(\theta)}f(x,\theta)$
    $=\left\{x\in C(\theta)\mid f(x,\theta)=f^{\star}(\theta)\right\}$
    is single-valued and continuous.
[…]
Proof.
As pointed out by Debreu, (1952), in finite dimension, Begle’s fixed point theorem applies to compact metric spaces that are both contractible and locally contractible. Local contractibility is a property of the space’s local geometry that holds for any subset of Euclidean space (see notes by Wilkins, (2017)). Hence, it suffices to assume contractibility. The statement then follows from Begle, (1950), Theorem 1. Note that we state the fact for single-valued functions which directly follows from the set-valued version. ∎
Appendix D Proof of Existence Result
In this section, we provide the formal proof of Theorem 3.2, based on the background introduced in Appendix C. First, we show playerwise convexity of the feasible slices.
Lemma 1.
Suppose Assumption 1 holds, and let $\mathcal{C}^{\prime}\subseteq\mathcal{C}$ be a connected component of $\mathcal{C}$. For any $i\in[m]$ and $x_{-i}\in\mathcal{X}_{-i}$, the set
$\mathcal{C}^{\prime}_{i}(x_{-i})\coloneqq\{x_{i}\in\mathcal{X}_{i}\mid(x_{i},x_{-i})\in\mathcal{C}^{\prime}\}$
is convex.
Proof.
If $\mathcal{C}^{\prime}_{i}(x_{-i})=\emptyset$, then convexity trivially holds. Otherwise, let $x_{i},x_{i}^{\prime}\in\mathcal{C}^{\prime}_{i}(x_{-i})$ be arbitrary. We know that $c_{j}(x_{i},x_{-i})\geq\alpha_{j}$ and $c_{j}(x_{i}^{\prime},x_{-i})\geq\alpha_{j}$ for all $j\in[b]$. By concavity of each $c_{j}$, we also have $c_{j}(x_{i}^{\lambda},x_{-i})\geq\alpha_{j}$ for any $x_{i}^{\lambda}\coloneqq\lambda x_{i}+(1-\lambda)x_{i}^{\prime}$ with $\lambda\in[0,1]$. Hence $(x_{i}^{\lambda},x_{-i})\in\mathcal{C}$. Moreover, $(x_{i}^{\lambda},x_{-i})$ must be in the same connected component of $\mathcal{C}$ as $(x_{i},x_{-i})$, otherwise we get a contradiction with the fact that the line segment $(x_{i},x_{-i}),(x_{i}^{\lambda},x_{-i})]\subset\mathcal{C}$. Therefore, we conclude that $(x_{i}^{\lambda},x_{-i})\in\mathcal{C}^{\prime}$ which proves convexity of $\mathcal{C}^{\prime}_{i}(x_{-i})$. ∎
Let $\operatorname{supp}(\mathcal{C}^{\prime}_{i})\coloneqq\left\{x_{-i}\in\mathcal{X}_{-i}\mid\exists x_{i}\in\mathcal{X}_{i}\text{ s.t.\ }(x_{i},x_{-i})\in\mathcal{C}^{\prime}\right\}$. We show continuity of $\mathcal{C}_{i}^{\prime}$ over this domain as follows.
Lemma 2.
Suppose Assumption 2 holds, and let $\mathcal{C}^{\prime}\subseteq\mathcal{C}$ be a connected component of $\mathcal{C}$. Then, for all $i\in[m]$, the set-valued function $\mathcal{C}^{\prime}_{i}(\cdot)$ is continuous (i.e. both upper and lower semicontinuous) over $\operatorname{supp}(\mathcal{C}^{\prime}_{i})$.
Proof.
We apply Lemma 6.1 of Still, (2018)’s notes on parametric optimization. In our setting, it states that if the MFCQ condition holds at every $x\in\mathcal{C}^{\prime}$ (which is given by our Assumption 2), then the parametric set-valued function $\mathcal{C}^{\prime}_{i}(\cdot)$ is $L$-Lipschitz continuous in terms of the Hausdorff distance over $\operatorname{supp}(\mathcal{C}^{\prime}_{i})$ for large enough $L>0$. It is well-known that for closed-valued functions such Hausdorff continuity implies upper and lower semicontinuity (see e.g. Clarke et al., (1998), 7.29). ∎
[…]
Proof.
Let $\mathcal{C}^{\prime}\subseteq\mathcal{C}$ be any connected component of $\mathcal{C}$, which exists as $\mathcal{C}\not$$\emptyset$. The proof is divided into two parts:
  1. 1)
    Contractibility of $\mathcal{C}^{\prime}$: For a compact set $X\subset\mathbb{R}^{n}$, its centroid is defined as
    $\overline{c}(X)\coloneqq\frac{1}{\text{vol}(X)}\int_{X}x\;dx$
    where $\text{vol}(X)$ is the $n$-dimensional volume of $X$ which is assumed to be positive [^1] Note that if $X$ collapses to a dimension lower than $n$, the centroid remains well-defined by using the volume of the respective dimension. If $X$ is a single point, then $\overline{c}(X)$ is that point..
    Towards constructing a contraction map, for $i\in[m]$, let $h_{i}:[0,1]\times\mathcal{C}^{\prime}\to\mathcal{C}^{\prime}$ be defined as
    $h_{i}(t,x)\coloneqq(1-t)x+t\left(\overline{c}(\mathcal{C}^{\prime}_{i}(x_{-i})),x_{-i}\right),$
    which can be seen as a mapping that contracts $\mathcal{C}^{\prime}$ along player $i$’s dimension. Next, we want to sequentially combine these functions in a way that for $t\in[\frac{i-1}{m},\frac{i}{m}]$, the playerwise contraction is applied according to $h_{i}$.
    Let $\varphi_{i}(t):[\frac{i-1}{m},\frac{i}{m}]\to[0,1]$ be defined as $\varphi_{i}(t)\coloneqq mt-i+1$ which will be used as an auxiliary rescaling function. Its inverse is well-defined and will be denoted by $\varphi^{-1}_{i}(t):[0,1]\to[\frac{i-1}{m},\frac{i}{m}]$. Then, for $i\in[m]$, we define $H_{i}:(\frac{i-1}{m},\frac{i}{m}]\times\mathcal{C}^{\prime}\to\mathcal{C}^{\prime}$ as
    $H_{1}(t,x)$ $\coloneqq h_{1}(\varphi_{1}(t),x)$
    $H_{i}(t,x)$ $\coloneqq h_{i}\left(\varphi_{i}(t),H_{i-1}\left(\varphi_{i-1}^{-1}(1),x\right)\right)$
    Finally, our contraction map $H:[0,1]\times\mathcal{C}^{\prime}\to\mathcal{C}^{\prime}$ is defined as
    $H(t,x)\coloneqq\begin{dcases}x,&\text{ if }t=0;\\ H_{i}(t,x),&\text{ if }t\in(\textstyle\frac{i-1}{m},\textstyle\frac{i}{m}]\text{ for }i\in[m].\end{dcases}$
    To prove that $H$ is indeed a contraction, it remains to show:
    1. (a)
      Well-definedness: For all $t\in[0,1],x\in\mathcal{C}^{\prime}$, $H(t,x)\in\mathcal{C}^{\prime}$.
    2. (b)
      Continuity: $H$ is continuous in $(t,x)$.
    3. (c)
      Contraction to point: There exists $\hat{x}\in\mathcal{C}^{\prime}$ such that for all $x\in\mathcal{C}^{\prime}$, $H(0,x)=x$ and $H(1,x)=\hat{x}$.
    Regarding 1)(a), observe that each $h_{i}$ is well-defined: By definition, we have $x_{i}\in\mathcal{C}^{\prime}_{i}(x_{-i})$. By convexity of $\mathcal{C}^{\prime}_{i}(x_{-i})$, as shown in Lemma 1, it holds that $\overline{c}(\mathcal{C}^{\prime}_{i}(x_{-i}))\in\mathcal{C}^{\prime}_{i}(x_{-i})$. Thus for all $t\in[0,1]$, again by convexity, $(1-t)x_{i}+t\,\overline{c}(\mathcal{C}^{\prime}_{i}(x_{-i}))\in\mathcal{C}^{\prime}$. As $h_{i}$ keeps $x_{-i}$ fixed, we have $h_{i}(t,x)\in\mathcal{C}^{\prime}$ for all $t\in[0,1]$ and $x\in\mathcal{C}^{\prime}$. Then, by definition, $H_{1}$ is well-defined, and the same follows for each $H_{i}$ inductively. We conclude that $H$ is well-defined, as it is piecewise equivalent to the $H_{i}$’s.
    To show 1)(b), note that by Lemma 2, $\mathcal{C}^{\prime}_{i}(x_{-i})$ is continuous in $x_{-i}$. Then, continuity of each $h_{i}$ in $(t,x)$ follows together with the continuity of the centroid function $\overline{c}(\cdot)$. Similarly as in (a), the recursive definitions of each $H_{i}$ and the piecewise definition of $H$ lets us conclude that $H$ is indeed continuous in $(t,x)$. In particular, we point out that continuity at the boundary of the pieces is ensured, as for all $i\in[m]\setminus\left\{1\right\}$ and all $x\in\mathcal{C}^{\prime}$, by definition, $\lim_{t\to{(i-1)/m}^{+}}H_{i}(t,x)=H_{i-1}(\varphi^{-1}_{i-1}(1),x)$.
    For 1)(c), by definition we have $H(0,x)=x$ for all $x\in\mathcal{C}$. To argue about $H(1,\cdot)$, let $\mathcal{C}^{\prime(0)}\coloneqq\mathcal{C}^{\prime}$ and, for $i\in[m]$, let $\mathcal{C}^{\prime(i)}\subseteq\mathcal{C}^{\prime}$ denote the image of the function $x\mapsto H_{i}(1,x)$. Thus, $\mathcal{C}^{\prime(i)}\coloneqq\{H_{i}(\varphi^{-1}_{i}(1),x)\mid x\in\mathcal{C}^{\prime}\}$ which due to the recursive definition of $H_{i}$, for all $i\in[m]$, is equal to $\{h_{i}(1,x)\mid x\in\mathcal{C}^{\prime(i-1)}\}$. Let $d_{i}\coloneqq\text{dim}(\mathcal{X}_{i})$ and notice that $\text{dim}(\mathcal{C}^{\prime(0)})\leq\sum_{i\in[m]}d_{i}$. Next, we have $\mathcal{C}^{\prime(1)}=\{(\overline{c}(\mathcal{C}_{1}^{\prime}(x_{-i})),x_{-i})\mid\exists x_{1}\in\mathcal{X}_{1}\text{ s.t.}\;(x_{1},x_{-1})\in\mathcal{C}^{\prime(0)}\}$, meaning $\mathcal{C}^{\prime(1)}$ is can be described as a parameterized set with parameters from a set of dimension at most $\sum_{i=2}^{m}d_{i}$. Hence $\text{dim}(\mathcal{C}^{\prime(1)})\leq\sum_{i=2}^{m}d_{i}$. Similarly, we can argue that for all $1<i\leq[m]$, $\text{dim}(\mathcal{C}^{\prime(i)})\leq\sum_{j=i+1}^{m}d_{i}$, since $\mathcal{C}^{\prime(i)}$ can be described as a parameterized set where coordinates corresponding to the first $i$ players are determined as a function of the remaining players $i+1$ to $m$. In particular, for the image of $x\mapsto H(1,x)$, namely the set $\mathcal{C}^{\prime(m)}$, we have therefore have $\text{dim}(\mathcal{C}^{\prime(m)})\leq 0$. Finally, in order to argue that $\mathcal{C}^{\prime(m)}$ is connected, we recall that the image of any continuous function with connected domain is connected. Since $\mathcal{C}^{\prime}$ was chosen as a connected component, by continuity of $x\mapsto H(1,x)$, $\mathcal{C}^{\prime(m)}$ is therefore connected. As a connected set of dimension 0, $\mathcal{C}^{\prime(m)}$ can only contain a single point $\hat{x}$, meaning $H(1,x)=\hat{x}$ for all $x\in\mathcal{C}$.
  2. 2)
    Fixed point argument: For any $\epsilon>0$, define the perturbed game $\Gamma^{\epsilon}$ with regularized utilities $u_{i}^{\epsilon}(x)\coloneqq u_{i}(x)-\epsilon\lVert x_{i}\rVert^{2}$; constraints are left unchanged. Observe that $\Gamma^{\epsilon}$ has playerwise strongly concave utilities. We first show that for any $\epsilon>0$, $\Gamma^{\epsilon}$ has a constrained Nash equilibrium in $\mathcal{C}^{\prime}$. Then we obtain the existence result for the original game via a limit argument.
    Let $\epsilon>0$. For each $i\in[m]$, define the playerwise best response map $\Psi_{i}:\mathcal{C}^{\prime}\to\mathcal{C}^{\prime}$ as
    $\Psi_{i}(x)\coloneqq\Big(\operatorname*{arg\,max}_{x_{i}^{\prime}\in\mathcal{C}^{\prime}_{i}(x_{-i})}u^{\epsilon}_{i}(x_{i}^{\prime},x_{-i}),\;x_{-i}\Big).$
    By a variant of Berge’s maximum theorem for strongly concave objective functions over convex constraints (see Theorem 1 in Appendix C), $\Psi_{i}$ is continuous and single-valued. Define the joint best response map $\Psi:\mathcal{C}^{\prime}\to\mathcal{C}^{\prime}$ as the composition $\Psi\coloneqq\Psi_{1}\circ\dots\circ\Psi_{m}$, which is also continuous.
    Since $\mathcal{C}^{\prime}$ is compact and contractible, by Fact 3.2, there exists $x^{\star}\in\mathcal{C}^{\prime}$ such that $x^{\star}=\Psi(x^{\star})$. As each $\Psi_{i}$ leaves the strategies of players other than $i$ unchanged, this implies that $\Psi_{i}(x^{\star})=x^{\star}$ for all $i\in[m]$. Suppose $x^{\star}$ is not a constrained Nash equilibrium. Then for some $i\in[m]$, there exists $x_{i}^{\prime}$ such that $\hat{x}\coloneqq(x_{i}^{\prime},x^{\star}_{-i})\in\mathcal{C}$ and $u_{i}(\hat{x})>u_{i}(x^{\star})$. We observe that by playerwise concavity of constraints, the line segment $x^{\star},\hat{x}]\subset\mathcal{C}$ satisfies all constraints, meaning $x^{\star}$ and $\hat{x}$ lie in the same connected component of $\mathcal{C}$; thus $\hat{x}\in\mathcal{C}^{\prime}\subseteq\mathcal{C}$. We have reached a contradiction with $x^{\star}_{i}$ being the maximizer in the definition of $\Psi_{i}(x^{\star})$, hence $x^{\star}$ is a constrained Nash equilibrium of $\Gamma^{\epsilon}$.
    Let $\smash{\epsilon^{(n)}}$ be a sequence so that $\smash{\epsilon^{(n)}}>0$ and $\smash{\epsilon^{(n)}}\to 0$. Let $\smash{x^{(n)}_{\star}}\in\mathcal{C}^{\prime}$ be a constrained Nash equilibrium of $\smash{\Gamma^{\epsilon^{(n)}}}$; this was just shown to exist. Let $x^{\star}$ be a limit point of the sequence $\smash{x^{(n)}_{\star}}$, which exists as $\smash{x^{(n)}_{\star}}\in\mathcal{C}^{\prime}$ and $\mathcal{C}^{\prime}$ is compact. For any $i\in[m]$, let $x_{i}^{\prime}\in\mathcal{C}_{i}(x^{\star}_{-i})$. Again, by playerwise concavity of constraints, $(x_{i}^{\prime},x^{\star}_{-i})\in\mathcal{C}^{\prime}$. By continuity of $u_{i}$, and using the fact that $\smash{x^{(n)}_{\star}}$ is a constrained Nash equilibrium of the perturbed game $\smash{\Gamma^{\epsilon^{(n)}}}$, we have
    $u_{i}(x_{i}^{\prime},x^{\star})=\lim_{n\to\infty}u_{i}^{\epsilon^{(n)}}(x_{i}^{\prime},x^{(n)}_{\star})\leq\lim_{n\to\infty}u_{i}^{\epsilon^{(n)}}(x^{(n)}_{\star})=u_{i}(x^{\star}),$
    i.e., we have shown that $x^{\star}$ is a constrained Nash equilibrium of the original unperturbed game.
Appendix E Background for Section 4
In this section, we introduce some background on constrained optimization relevant for the convergence analyses in Appendix F. First, we establish connections between KKT conditions and constrained Nash equilibria. Second, we prove a technical lemma which will be applied to establish a lower bound on the adaptive stepsizes.
E.1 Lagrangian, KKT conditions, and connections to constrained Nash equilibria
We define Lagrangian functions in a playerwise manner, i.e., with respect to a fixed strategy for all except one player. Formally, for $i\in[m]$ and $x_{-i}\in\mathcal{X}_{-i}$, we define $\mathcal{L}_{i}^{x_{-i}}:\mathcal{X}_{i}\times\mathbb{R}^{b}\to\mathbb{R}$ as
$\mathcal{L}_{i}^{x_{-i}}(x_{i},\lambda)\coloneqq u_{i}(x_{i},x_{-i})+\sum_{j\in[b]}\lambda_{j}\left(c_{j}(x_{i},x_{-i})-\alpha_{j}\right).$
Next, we introduce approximate KKT conditions that are formulated based on the above Lagrangian in a playerwise manner.
Definition 2.
For $\epsilon>0$, a strategy profile $x\in\mathcal{X}$ satisfies our $\epsilon$-KKT conditions, if there exist Lagrange multipliers $\lambda=\left(\lambda_{1},\dots,\lambda_{b}\right)\in\mathbb{R}^{b}$ such that
$x\in\mathcal{C}\hfill$ (primal feasibility)
$\min_{j\in[b]}\lambda_{j}\geq 0\hfill$ (dual feasibility)
$\max_{j\in[b]}\left|\lambda_{j}\left(c_{j}(x)-\alpha_{j}\right)\right|\leq\epsilon\hfill$ (complementary slackness)
$\forall i\in[m],\quad\max_{x^{\prime}_{i}\in\mathcal{X}_{i}}\langle x_{i}^{\prime}-x_{i},\nabla_{x_{i}}\mathcal{L}^{x_{-i}}_{i}(x_{i},\lambda)\rangle\leq\epsilon\hfill$ (Lagrangian stationarity)
Next, we establish the following connection between these approximate KKT conditions and approximate constrained Nash equilibria.
Lemma 3.
Let $\epsilon>0$ and let $(x,\lambda)\in\mathcal{X}\times\mathbb{R}^{b}$ satisfy the above $\epsilon$-KKT conditions. Then $x$ is a constrained $(2\epsilon)$-approximate Nash equilibrium.
Proof.
Let $i\in[m]$ and $x_{i}^{\prime}\in\mathcal{C}_{i}(x_{-i})$ be any feasible deviation of player $i$. Then, using concavity of $u_{i}$ and the definition of the Lagrangian, we bound
$u_{i}(x_{i}^{\prime},x_{-i})-u_{i}(x_{i},x_{-i})$ $\leq\left\langle x_{i}^{\prime}-x_{i},\nabla_{x_{i}}u_{i}(x_{i},x_{-i})\right\rangle$ (3)
$=\left\langle x^{\prime}_{i}-x_{i},\nabla_{x_{i}}\mathcal{L}_{i}^{x_{-i}}(x_{i})\right\rangle-\sum_{j\in[b]}\lambda_{j}\left\langle x^{\prime}_{i}-x_{i},\nabla_{x_{i}}c_{j}(x_{i},x_{-i})\right\rangle.$
For each $j\in[b]$, we can bound the respective term in the sum on the right-hand side using concavity of $c_{j}$ by
$\lambda_{j}\left\langle x^{\prime}_{i}-x_{i},\nabla_{x_{i}}c_{j}(x_{i},x_{-i})\right\rangle$ $\geq\lambda_{j}c_{j}(x^{\prime}_{i},x_{-i})-\lambda_{j}c_{j}(x_{i},x_{-i})$ (4)
$=\lambda_{j}\left(c_{j}(x^{\prime}_{i},x_{-i})-\alpha_{j}\right)-\lambda_{j}\left(c_{j}(x_{i},x_{-i})-\alpha_{j}\right)$
$\geq-\lambda_{j}\left(c_{j}(x_{i},x_{-i})-\alpha_{j}\right)$
where the last step is due to $x_{i}^{\prime}\in\mathcal{C}_{i}(x_{-i})$, i.e., $c_{j}(x^{\prime}_{i},x_{-i})\geq\alpha_{j}$. Plugging (4) into (3), and applying Lagrangian stationarity and complementary slackness, we get
$u_{i}(x_{i}^{\prime},x_{-i})-u_{i}(x_{i},x_{-i})\leq\left\langle x^{\prime}_{i}-x_{i},\nabla_{x_{i}}\mathcal{L}_{i}^{x_{-i}}(x_{i})\right\rangle+\sum_{j\in[b]}\lambda_{j}\left(c_{j}(x_{i},x_{-i})-\alpha_{j}\right)\leq 2\epsilon$
which implies the claimed statement by definition of approximate constrained Nash equilibrium. ∎
E.2 Technical lemma
Notation
For $x,x^{\prime}\in\mathbb{R}^{n}$, we denote the line segement connecting $x$ and $x^{\prime}$ by
$[x,x^{\prime}]\coloneqq\left\{\lambda x+(1-\lambda)x^{\prime}\mid\lambda\in[0,1]\right\}.$
We provide a lemma which is applied in the proof of Lemma 11 in Appendix F. Informally, it says the following: If we perform a projected gradient ascent step along a (locally) smooth function $f$ and the observed gradients of $f$ are perturbed by some error $\delta$, then, as long as $\delta$ is sufficiently small, the value of $f$ will not decrease.
Lemma 4.
Let $\Omega\subset\mathbb{R}^{n}$ and let $f:\Omega\to\mathbb{R}$ be differentiable. For $\kappa\geq 1$, let
$x^{\prime}$ $\coloneqq x+\frac{1}{\kappa}\nabla f(x),$
$x^{+}$ $\coloneqq\mathcal{P}_{\Omega}\left[x^{\prime}\right],$
and define the mapping $F(x)\coloneqq\kappa(x^{+}-x)$. Suppose for all $\hat{x}\in[x,x^{+}]$, it holds that
$\left\lVert\nabla f(x^{+})-\nabla f(x)\right\rVert\leq\kappa\lVert x^{+}-x\rVert.$
For $\delta\in\mathbb{R}^{n}$, define the perturbed update as
$\widetilde{x}^{\prime}$ $\coloneqq x+\frac{1}{\kappa}\left(\nabla f(x)+\delta\right),$
$\widetilde{x}^{+}$ $\coloneqq\mathcal{P}_{\Omega}\left[\widetilde{x}^{\prime}\right$].
If $\lVert F(x)\rVert\geq 3\lVert\delta\rVert$, then $f(\widetilde{x}^{+})\geq f(x)$.
Proof.
[…]
  • Step 1: First, we show that $\lVert\widetilde{F}(x)\rVert\geq 2\lVert\delta\rVert$. Namely, we can bound
    $\lVert\widetilde{F}(x)\rVert$ $\overset{a)}{\geq}\lVert F(x)\rVert-\lVert\widetilde{F}(x)-F(x)\rVert$
    $\overset{b)}{\geq}3\lVert\delta\rVert-\kappa\lVert\widetilde{x}^{+}-x^{+}\rVert$
    $\overset{c)}{\geq}3\lVert\delta\rVert-\kappa\lVert\widetilde{x}^{\prime}-x^{\prime}\rVert$
    $\overset{d)}{=}3\lVert\delta\rVert-\lVert\delta\rVert$
    $=2\lVert\delta\rVert.$
    where a) is by the triangle inequality, b) by assumption of the lemma and definition of $\widetilde{F}$ and $F$, c) by non-expansiveness of projection, and d) by definition of the perturbed update.
  • Step 2: By the projection theorem, for all $z\in\Omega$, it holds that $(\widetilde{x}^{\prime}-\widetilde{x}^{+})^{\top}(z-\widetilde{x}^{+})\leq 0$. Thus, with $z=x$ and $\widetilde{x}^{+}-x=\frac{1}{\kappa}\widetilde{F}(x)$, we get
    $(\widetilde{x}^{\prime}-x+x-\widetilde{x}^{+})^{\top}(z-\widetilde{x}^{+})=\left(\frac{1}{\kappa}(\nabla f(x)+\delta)-\frac{1}{\kappa}\widetilde{F}(x)\right)^{\top}\left(-\frac{1}{\kappa}\widetilde{F}(x)\right)\leq 0.$
    After rearranging, this gives
    $\nabla f(x)^{\top}\widetilde{F}(x)\geq-\delta^{\top}\widetilde{F}(x)+\lVert\widetilde{F}(x)\rVert^{2}.$ (5)
    Then, starting from the $\kappa$-smoothness lower bound, we get
[…]
    $\overset{a)}{\geq}f(x)+\frac{1}{\kappa}\nabla f(x)^{\top}\widetilde{F}(x)-\frac{1}{2\kappa}\lVert\widetilde{F}(x)\rVert^{2}$
    $\overset{b)}{\geq}f(x)-\frac{1}{\kappa}\delta^{\top}\widetilde{F}(x)+\frac{1}{2\kappa}\lVert\widetilde{F}(x)\rVert^{2},$
    $\overset{c)}{\geq}f(x)-\frac{1}{\kappa}\lVert\delta\rVert\cdot\lVert\widetilde{F}(x)\rVert+\frac{1}{2\kappa}\lVert\widetilde{F}(x)\rVert^{2}$
    $\overset{d)}{\geq}f(x)-\frac{1}{2\kappa}\lVert\widetilde{F}(x)\rVert^{2}+\frac{1}{2\kappa}\lVert\widetilde{F}(x)\rVert^{2}$
    $\geq f(x)$
    where a) is by definition of $\widetilde{F}$, b) uses inequality (5), c) is by the Cauchy-Schwarz inequality, and d) by the bound shown in Step 1.
We restate the following sufficient increase inequality for projected gradient ascent that will be used in the proof of Theorem 3. The claimed inequality is standard; for reference, see Bubeck et al., (2015), Lemma 3.6.
Lemma 5.
Let $\Omega\subset\mathbb{R}^{n}$ and let $f:\Omega\to\mathbb{R}$ be differentiable. For $\kappa\geq 1$, let
$x^{+}\coloneqq\mathcal{P}_{\Omega}\left[x+\frac{1}{\kappa}\nabla f(x)\right],$
and suppose for all $\hat{x}\in[x,x^{+}]$,
$\lVert\nabla f(\hat{x})-\nabla f(x)\rVert\leq\kappa\lVert\hat{x}-x\rVert.$
Then, we have
$f(x^{+})-f(x)\geq\frac{\kappa}{2}\lVert x-x^{+}\rVert^{2}.$
Appendix F Proofs for Section 4.1
In this section, we provide the full proofs for the results of Section 4.1.
F.1 Mixed-strategy potential games
In Example 3 we claimed that when considering mixed strategies over a finite action potential game, the potential structure is lifted to the continuous game over compact convex set of mixed strategies. Here we state the corresponding proposition which is a direct consequence of Monderer and Shapley, (1996), Lemma 2.10.
Proposition 2.
Let $\mathcal{A}\coloneqq\times_{i\in[m]}\mathcal{A}_{i}$ with each $\mathcal{A}_{i}$ finite, and suppose $\widetilde{u}_{i}:\mathcal{A}\to[0,1]$ are player utilities such that there exists a potential function $\phi:\mathcal{A}\to\mathbb{R}$ satisfying
$\phi(a)-\phi(a_{i}^{\prime},a_{-i})=\widetilde{u}_{i}(a)-\widetilde{u}_{i}(a_{i}^{\prime},a_{-i})\quad\text{for all }i\in[m],\,a\in\mathcal{A},\,a_{i}^{\prime}\in\mathcal{A}_{i}.$
Then, the continuous game with strategy space $\mathcal{X}\coloneqq\times_{i\in[m]}\Delta(\mathcal{A}_{i})$ and expected utilities
$u_{i}(x)\coloneqq\mathbb{E}_{a\sim x}[\widetilde{u}_{i}(a)]$
is a potential game with potential function $\Phi(x)\coloneqq\mathbb{E}_{a\sim x}[\phi(a)]$.
F.2 Technical lemmas
Next, we provide several technical lemmas that will be used in our convergence analysis. First, we show that the log barrier regularized game is still a potential game.
Lemma 6.
Suppose $\Phi:\mathcal{X}\to\mathbb{R}$ is a potential function for the game with utilities $\{u_{i}\}_{i\in[m]}$. Then $\Phi^{\eta}(x)\coloneqq\Phi(x)+\eta\sum_{j\in[b]}\log(c_{j}(x)-\alpha_{j})$ is a potential for the log barrier regularized game with utilities $\{B^{\eta}_{i}\}_{i\in[m]}$.
Proof.
It is immediate from Assumption 6 that if we have a game with utilities $\{u^{\prime}_{i}\}_{i\in[m]}$ and potential function $\Phi^{\prime}$, and another game with utilities $\{u^{\prime\prime}_{i}\}_{i\in[m]}$ and potential function $\Phi^{\prime\prime}$, then $\Phi^{\prime}+\Phi^{\prime\prime}$ is a potential function for the game with utilities $\{u^{\prime}_{i}+u^{\prime\prime}_{i}\}_{i\in[m]}$. The statement hence follows by observing that the sum of log barrier terms forms an identical interest game and hence $\eta\sum_{j\in[b]}\log(c_{j}(x)-\alpha_{j})$ is a potential function for it. ∎
Next, we show that the potential function of the unregularized game is smooth.
Lemma 7.
Let $\Phi:\mathcal{X}\to\mathbb{R}$ be a potential function for our game with utilities $\{u_{i}\}_{i\in[m]}$. Then $\Phi$ is $(M\sqrt{m})$-smooth in the joint strategy $x\in\mathcal{X}$, where $M$ is the smoothness parameter of the utilities, see Assumption 3.
Proof.
Note that differentiability of $u_{i}$ for all $i\in[m]$ (Assumption 3) implies differentiability of $\Phi$, and it is well known that for all $x\in\mathcal{X}$, $\nabla_{x_{i}}\Phi(x)=\nabla_{x_{i}}u_{i}(x)$. Using the smoothness of $u_{i}$ from Assumption 3, we get
$\lVert\nabla\Phi(x)-\nabla\Phi(x^{\prime})\rVert^{2}$ $=\sum_{i\in[m]}\lVert\nabla_{x_{i}}\Phi(x)-\nabla_{x_{i}}\Phi(x^{\prime})\rVert^{2}$
$=\sum_{i\in[m]}\lVert\nabla_{x_{i}}u_{i}(x)-\nabla_{x_{i}}u_{i}(x^{\prime})\rVert^{2}$
$\leq M^{2}\sum_{i\in[m]}\lVert x-x^{\prime}\rVert^{2}$
$\leq M^{2}m\lVert x-x^{\prime}\rVert^{2}$
from which we conclude that $\lVert\nabla\Phi(x)-\nabla\Phi(x^{\prime})\rVert\leq M\sqrt{m}\lVert x-x^{\prime}\rVert$. ∎
F.3 Smoothness along the trajectory
In the following subsections, we present a convergence analysis of our independent log barrier method applied to constrained games with potential structure. We begin by examining the smoothness of properties of $\Phi^{\eta}$ along the joint gradient ascent trajectory. For this, we introduce the following notion of $\beta$-feasibility.
Definition 3.
A strategy $x\in\mathcal{C}$ is $\beta$-feasible for $\beta\geq 0$, if $c_{j}(x)-\alpha_{j}\geq\beta$ for all $j\in[b]$.
First, we show that $\beta$-feasibility can be preserved along the line segment spanned by consecutive iterates generated by update (1), here up to a factor of $1/2$, when choosing sufficiently small stepsizes.
Lemma 8.
Let $x^{(t)}\in\mathcal{C}$ be $\beta$-feasible. If $\gamma^{(t)}\leq\frac{\beta^{2}}{2mL^{2}(\beta+\eta b)}$, then for all $\hat{x}\in[x^{(t)},x^{(t+1)}]$, $\hat{x}$ is $\beta/2$-feasible.
Proof.
First, we use $\beta$-feasibility of $x^{(t)}$ to bound the norm of log barrier regularized partial gradients. Namely, for any $i\in[m]$, we have
$\lVert\nabla_{x_{i}}B_{i}^{\eta}(x^{(t)})\rVert\leq\lVert\nabla_{x_{i}}u_{i}(x^{(t)})\rVert+\eta\sum_{j\in[b]}\frac{\lVert\nabla_{x_{i}}c_{j}(x^{(t)})\rVert}{\beta}\leq\left(1+\frac{\eta b}{\beta}\right)L=\frac{(\beta+\eta b)L}{\beta}$
where we use $L$-Lipschitz continuity of $u_{i}$ and $c_{j}$, see Assumption 3, for bounding the gradient norms. Then, we have for each $j\in[b]$,
$c_{j}(\hat{x})-\alpha_{j}$ $\geq c_{j}(x^{(t)})-\alpha_{j}-L\lVert\hat{x}-x^{(t)}\rVert$
$\geq c_{j}(x^{(t)})-\alpha_{j}-L\lVert x^{(t+1)}-x^{(t)}\rVert$
$\geq c_{j}(x^{(t)})-\alpha_{j}-L\sum_{i\in[m]}\lVert x_{i}^{(t+1)}-x_{i}^{(t)}\rVert$
$\overset{a)}{\geq}c_{j}(x^{(t)})-\alpha_{j}-L\sum_{i\in[m]}\gamma^{(t)}\lVert\nabla_{x_{i}}B^{\eta}_{i}(x^{(t)})\rVert$
$\overset{b)}{\geq}c_{j}(x^{(t)})-\alpha_{j}-\frac{\beta}{2}$
$\overset{c)}{\geq}\frac{c_{j}(x^{(t)})-\alpha_{j}}{2}$
where a) is by non-expansiveness of the projection operator, b) by our choice of $\gamma^{(t)}$, and c) by $\beta$-feasibility of $x^{(t)}$. ∎
For $\beta$-feasible strategies, we give the following Hessian norm bound.
Lemma 9.
Let $x\in\mathcal{C}$ be $\beta$-feasible. Then $\lVert\nabla^{2}\Phi^{\eta}(x)\rVert\leq M^{\eta,\beta}_{\Phi}$ where
$M^{\eta,\beta}_{\Phi}\coloneqq M\sqrt{m}+\eta Mb\beta^{-1}+\eta L^{2}b\beta^{-2}.$
Proof.
For any $x\in\mathcal{X}$, we have
$\nabla^{2}\Phi^{\eta}(x)=\nabla^{2}\Phi(x)-\eta\sum_{j\in[b]}\frac{\nabla^{2}c_{j}(x)}{c_{j}(x)-\alpha_{j}}+\eta\sum_{j\in[b]}\frac{\nabla c_{j}(x)\nabla c_{j}(x)^{\top}}{(c_{j}(x)-\alpha_{j})^{2}}.$
Using $M$-smoothness and $L$-Lipschitz continuity of $c_{j}$ for all $j\in[b]$, as well as $(M\sqrt{m})$-smoothness of $\Phi$, as shown in Lemma 7, we then bound the Hessian norm as follows:
$\lVert\nabla^{2}\Phi^{\eta}(x)\rVert$ $\leq\lVert\nabla^{2}\Phi(x)\rVert+\eta\sum_{j\in[b]}\frac{\lVert\nabla^{2}c_{j}(x)\rVert}{c_{j}(x)-\alpha_{j}}+\eta\sum_{j\in[b]}\frac{\lVert\nabla c_{j}(x)\nabla c_{j}(x)^{\top}\rVert}{(c_{j}(x)-\alpha_{j})^{2}}$
$\leq M\sqrt{m}+\eta\sum_{j\in[b]}\frac{M}{c_{j}(x)-\alpha_{j}}+\eta\sum_{j\in[b]}\frac{L^{2}}{(c_{j}(x)-\alpha_{j})^{2}}$
$\leq M\sqrt{m}+\eta Mb\beta^{-1}+\eta L^{2}b\beta^{-2}$
$=M_{\Phi}^{\eta,\beta}.$
The Hessian norm bound from above is used to establish smoothness along the trajectory as follows.
Lemma 10.
Suppose $x^{(t)}\in\mathcal{C}$ is $\beta$-feasible and each $\gamma^{(t)}$ is bounded as in the statement of Lemma 8. Then $\Phi^{\eta}$ is $M^{\eta,\beta/2}_{\Phi}$-smooth over $x^{(t)},x^{(t+1)}]$ where $M^{\eta,\beta}_{\Phi}$ is defined as in Lemma 9, i.e. for all $x,x^{\prime}\in[x^{(t)},x^{(t+1)}]$,
$\left\lVert\nabla\Phi^{\eta}(x)-\nabla\Phi^{\eta}(x^{\prime})\right\rVert\leq M^{\eta,\beta/2}_{\Phi}\left\lVert x-x^{\prime}\right\rVert.$
Proof.
By the fundamental theorem of calculus,
$\nabla\Phi(x)-\nabla\Phi(x^{\prime})=\int_{0}^{1}\nabla^{2}\Phi(x^{\prime}+\tau(x-x^{\prime}))(x-x^{\prime})\,d\tau.$
Taking the norm, and using the fact that $x^{\prime}+t(x-x^{\prime})$ is $\beta/2$-feasible for all $t\in[0,1]$ as shown in Lemma 8, together with the Hessian norm bound from Lemma 9, we get
$\lVert\nabla\Phi^{\eta}(x)-\nabla\Phi^{\eta}(x^{\prime})\rVert$ $\leq\int_{0}^{1}\lVert\nabla^{2}\Phi^{\eta}(x^{\prime}+\tau(x-x^{\prime}))\rVert\cdot\lVert x-x^{\prime}\rVert\,d\tau$
$\leq\int_{0}^{1}M^{\eta,\beta/2}_{\Phi}\lVert x-x^{\prime}\rVert\,d\tau$
$\leq M^{\eta,\beta/2}_{\Phi}\lVert x-x^{\prime}\rVert.$
F.4 Lower bounding the stepsize
Motivated by the bound in Lemma 8 and smoothness along the trajectory as shown in Lemma 10, we choose the stepsize at iteration $t$ as
$\gamma^{(t)}\coloneqq\min\left\{\frac{{\beta^{(t)}}^{2}}{2mL^{2}(\beta^{(t)}+\eta b)},\frac{1}{M^{\eta,\beta^{(t)}/2}_{\Phi}}\right\}$ (6)
where $\beta^{(t)}\coloneqq\min_{j\in[b]}\left\{c_{j}(x^{(t)})-\alpha_{j}\right\}$, i.e. such that $x^{(t)}$ is $\beta^{(t)}$-feasible [^2] Note that stepsizes (6) differ from what is specified as (2) in the main part of the paper. This is a typo and will be fixed in a later version..
In this subsection, we want to argue that the above defined stepsizes can be uniformly lower bounded by a constant multiple of $\eta$. This is necessary to guarantee sufficient progress in our convergence analysis. We start by showing that feasibility margins do not decrease further once a point sufficiently close to the boundary is reached. Based on Lemma 4, we prove the following statement. Note that a similar statement was proven in Usmanova et al., (2024), Fact 1, with the difference that ours is for projected gradient steps, which requires a different approach.
In the following, for $j\in[b]$, we write $\beta^{(t)}_{j}\coloneqq c_{j}(x^{(t)})-\alpha_{j}$.
Lemma 11.
Let Assumption 5 hold with $\rho\geq\eta$. If at any iteration $0\leq t\leq T-1$, we have $\beta^{(t)}\leq\overline{c}\eta$ with $\overline{c}=\frac{\ell}{6L(b+1)}$, then for iteration $t+1$, we get $\prod_{j\in\mathcal{J}}\beta_{j}^{(t+1)}\geq\prod_{j\in\mathcal{J}}\beta_{j}^{(t)}$ for any $\mathcal{J}$ such that $\mathcal{J}_{t}\subseteq\mathcal{J}$ with $\mathcal{J}_{t}\coloneqq\{j\in[b]\mid\beta_{j}^{(t)}\leq\eta\}$.
Proof.
We aim to apply Lemma 4 with $x^{(t)}, x+=x$$^{+}=x^{(t+1)}$,
$\delta$ $=\nabla\Phi(x^{(t)})+\eta\sum_{j\in[m]\setminus\mathcal{J}_{t}}\nabla\log\beta_{j}^{(t)},\text{ and}$
$f(x)$ $=\eta\sum_{j\in\mathcal{J}_{t}}\log\beta_{j}^{(t)}.$
Note that by our choice of $\gamma^{(t)}$, by Lemma 10, we have $\lVert\nabla f(x^{+})-\nabla f(x)\rVert\leq\frac{1}{\gamma^{(t)}}\lVert x^{+}-x\rVert$. In order to apply Lemma 4, we further need to show that for $F(x)\coloneqq\frac{1}{\gamma^{(t)}}\left(x^{+}-x\right)$, it holds that $\lVert F(x)\rVert\geq 3\lVert\delta\rVert$.
We begin by giving a lower bound on $\lVert F(x^{(t)})\rVert$. By Proposition B.1 of Agarwal et al., (2021) and Assumption 5, we have
$2\left\lVert F(x^{(t)})\right\rVert$ $\geq\max_{x^{(t)}+d\,\in\,\mathcal{X},\lVert d\rVert\leq 1}d^{\top}\nabla f(x^{(t)})$
$\geq\eta\max_{x^{(t)}+d\,\in\,\mathcal{X},\lVert d\rVert\leq 1}\sum_{j\in\mathcal{J}_{t}}\frac{d^{\top}\nabla c_{j}(x^{(t)})}{\beta_{j}^{(t)}}$
$\geq\eta\sum_{j\in\mathcal{J}_{t}}\frac{\ell}{\beta_{j}^{(t)}}$
$\geq\frac{\ell}{\overline{c}}.$
Then, we show an upper bound on $\lVert\delta\rVert$,
$\lVert\delta\rVert$ $=\left\lVert\nabla\Phi(x^{(t)})+\eta\sum_{j\in[m]\setminus\mathcal{J}_{t}}\nabla\log\beta_{j}^{(t)}\right\rVert$
$\leq\left\lVert\nabla\Phi(x^{(t)})\right\rVert+\eta\sum_{j\in[m]\setminus\mathcal{J}_{t}}\left\lVert\frac{\nabla c_{j}(x^{(t)})}{\beta_{j}^{(t)}}\right\rVert$
$\overset{(\star)}{\leq}L+\eta\cdot b\cdot\frac{L}{\eta}$
$\leq L(b+1),$
where $(\star)$ is due to $\beta_{j}^{(t)}>\eta$ for $j\not\in\mathcal{J}_{t}$. Combining the two bounds, and by definition of $\overline{c}$, we obtain from Lemma 4 that
$\eta\sum_{j\in\mathcal{J}_{t}}\log\beta^{(t+1)}_{j}\geq\eta\sum_{j\in\mathcal{J}_{t}}\log\beta^{(t)}_{j}$.
Moreover, using the same reasoning, we can prove this inequality to hold for the summation over $\mathcal{J}$ instead of $\mathcal{J}_{t}$. Then, dividing by $\eta$ and exponentiating both sides yields the claimed result. ∎
As a consequence of Lemma 11, we obtain the following lower bound on the distance of iterates to the boundary of $\mathcal{C}$. Having established Lemma 11 above, the proof of Lemma 12 is equivalent to the proof of an analogous result stated as Lemma 6 in Usmanova et al., (2024); we omit the proof here.
Lemma 12.
Let Assumptions 4 and 5 hold with $\rho\geq\eta$. Then, for all $t\geq 1$, it holds that $\beta^{(t)}\geq c\eta$ where $c\coloneqq\frac{1}{2}\left(\frac{\ell}{24L(b+1)}\right)^{b}$.
The above lower bound also implies a lower bound on the stepsizes.
Lemma 13.
If $\beta^{(t)}\geq c\eta$ for some $c>0$, then we have $\gamma^{(t)}\geq C\eta$ with
$C\coloneqq\min\left\{\frac{c}{2mL^{2}(1+b)},\frac{1}{M\sqrt{m}+2Mbc^{-1}+4L^{2}bc^{-2}}\right\}$
Proof.
Recall that we have set
$\gamma^{(t)}\coloneqq\min\left\{\frac{{\beta^{(t)}}^{2}}{2mL^{2}(\beta^{(t)}+\eta b)},\frac{1}{M^{\eta,\beta^{(t)}/2}_{\Phi}}\right\}.$
As we are interested in small $\eta>0$, we may assume w.l.o.g. that $\eta\leq 1$. We may further assume that $\beta^{(t)}\leq 1$; otherwise, we trivially have a lower bound on $\beta^{(t)}$. Then the first argument of the minimum above is lower bounded by
$\frac{{\beta^{(t)}}^{2}}{2mL^{2}(\beta^{(t)}+\eta b)}\geq\frac{\beta^{(t)}}{2mL^{2}(1+\eta b)}\geq\frac{\eta c}{2mL^{2}(1+b)}$
For the second term, by definition of $M^{\eta,\beta}_{\Phi}$, we get
$\frac{1}{M^{\eta,\beta^{(t)}/2}_{\Phi}}$ $\geq\frac{1}{M\sqrt{m}+\frac{2Mb}{c}+\frac{4L^{2}b}{c^{2}\eta}}$
$\geq\frac{\eta}{\eta M\sqrt{m}+\frac{2\eta Mb}{c}+\frac{4L^{2}b}{c^{2}}}$
$\geq\frac{\eta}{M\sqrt{m}+\frac{2Mb}{c}+\frac{4L^{2}b}{c^{2}}}$
The claimed bound thus follows by combining both. ∎
Finally, we are ready to prove the main convergence result for the constrained potential game setting.
Theorem 3.
Let $\epsilon>0$ and set $\eta=\epsilon$. Let Assumptions 1, 3, 4, 6, and 5 hold with $\rho\geq\eta$. For stepsizes $\{\gamma^{(t)}\}_{t=0}^{T-1}$ chosen as in (6), suppose we perform $T=\mathcal{O}(\epsilon^{-3})$ rounds of simultaneous updates as in (1). Then, there exists $0\leq t\leq T$ such that $x^{(t)}$ is a constrained $\epsilon$-approximate Nash equilibrium.
Proof.
First, we argue that our independent updates are equivalent to centralized gradient ascent on the regularized potential function $\Phi^{\eta}$. This follows from the fact that, by Lemma 6, the regularized utilities form a potential game, meaning we have $\nabla_{x_{i}}\Phi^{\eta}(x)=\nabla_{x_{i}}B^{\eta}_{i}(x)$ for any $x\in\mathcal{X}$ and $i\in[m]$. Together with separability of the projection operator onto the Cartesian product of the sets $\mathcal{X}_{i}$, see Leonardos et al., (2022), Lemma D.1, we obtain
$\left(\mathcal{P}_{\mathcal{X}_{i}}\left[x^{(t)}_{i}+\gamma^{(t)}\nabla_{x_{i}}B^{\eta}_{i}(x^{(t)})\right]\right)_{i\in[m]}=\mathcal{P}_{\mathcal{X}}\left[x^{(t)}+\gamma^{(t)}\nabla\Phi^{\eta}(x^{(t)})\right].$
Hence, in the following we carry out the analysis by viewing the algorithm as a centralized gradient ascent on $\Phi^{\eta}$.
By smoothness along the trajectory and our choice of stepsizes, by Lemma 5, for all $0\leq t\leq T-1$,
$\Phi^{\eta}(x^{(t+1)})-\Phi^{\eta}(x^{(t)})$ $\geq\frac{\gamma^{(t)}}{2}\left\lVert x^{(t+1)}-x^{(t)}\right\rVert^{2}.$
Together with the fact that there exists a constant $C>0$ such that $\gamma^{(t)}\geq C\eta$, see Lemma 13, and by telescoping the sum, we obtain
$\frac{1}{T}\sum_{t=0}^{T-1}\left\lVert x^{(t+1)}-x^{(t)}\right\rVert^{2}$ $\leq\frac{2}{T}\sum_{t=0}^{T-1}\frac{\Phi^{\eta}(x^{(t+1)})-\Phi^{\eta}(x^{(t)})}{\gamma^{(t)}}$
$\leq\frac{2\Phi^{\eta}(x^{(T)})}{C\eta T}.$
We note that using Lemma 12, $\Phi^{\eta}$ can be bounded by a constant. Thus for $\eta=\epsilon$ and $T=\mathcal{O}(\epsilon^{-3})$, there exists $0\leq t^{\star}\leq T-1$ such that $\lVert x^{(t^{\star}+1)}-x^{(t^{\star})}\rVert\leq\frac{\epsilon}{4}$. Next, we argue that then $x^{(t^{\star}+1)}$ satisfies the $\frac{\epsilon}{2}$-KKT conditions as introduced in Appendix E. Consider the pair of primal and dual variables given by
$(x^{\star},\lambda^{\star})\coloneqq\left(x^{(t^{\star}+1)},\;\frac{\eta}{2}\left[(c_{j}(x^{(t^{\star}+1)})-\alpha_{j})^{-1}\right]_{j\in[b]}\right).$
We check each of the conditions:
  1. 1.
    Primal feasibility: Lemma 8 implies that all iterates remain within $\mathcal{C}$. In particular, we have $x^{\star}\in\mathcal{C}$.
  2. 2.
    Dual feasibility: This follows directly from primal feasibility which implies that $c_{j}(x^{\star})-\alpha_{j}\geq 0$ for all $j\in[b]$, and thus $\lambda^{\star}_{j}\geq 0$.
  3. 3.
    Complementary slackness: This is a consequence of the definition of $\lambda^{\star}$, as
    $\max_{j\in[b]}\left|\lambda^{\star}_{j}(c_{j}(x^{\star})-\alpha_{j})\right|=\frac{\eta}{2}=\frac{\epsilon}{2}.$
  4. 4.
    Lagrangian stationarity: By a standard property of projected gradient ascent, see e.g. Agarwal et al., (2021), Proposition B.1, $\lVert x^{(t^{\star}+1)}-x^{(t^{\star})}\rVert\leq\frac{\epsilon}{4}$ implies that
    $\max_{x\in\mathcal{X}}\left\langle x-x^{\star},\nabla\Phi^{\eta}(x^{\star})\right\rangle\leq\frac{\epsilon}{2}.$
    Then we can bound, for any $i\in[m]$,
    $\max_{x_{i}\in\mathcal{X}_{i}}\left\langle x^{\star}_{i}-x_{i},\nabla_{x_{i}}\mathcal{L}_{i}^{x^{\star}_{-i}}(x^{\star}_{i},\lambda^{\star})\right\rangle$ $=\max_{x_{i}\in\mathcal{X}_{i}}\left\langle x^{\star}-(x_{i},x^{\star}_{-i}),\nabla\Phi^{\eta}(x^{\star})\right\rangle$
    $\leq\max_{x\in\mathcal{X}}\left\langle x^{\star}-x,\nabla\Phi^{\eta}(x^{\star})\right\rangle$
    $\leq\frac{\epsilon}{2},$
    where the first step is by definition of $\lambda^{\star}$.
Finally, Lemma 3 implies that the $\frac{\epsilon}{2}$-KKT strategy $x^{\star}$ constitutes an $\epsilon$-approximate constrained Nash equilibrium. ∎